### v3

# Agentic Large Language Models, a survey

Abstract.
Background: There is great interest in agentic LLMs, large language models that act as agents.
Objectives: We review the growing body of work in this area and provide a research agenda.
Methods: Agentic LLMs are LLMs that (1) reason, (2) act, and (3) interact. We organize the literature according to these three categories.
Results: The research in the first category focuses on reasoning, reflection, and retrieval, aiming to improve decision making; the second category focuses on action models, robots, and tools, aiming for agents that act as useful assistants; the third category focuses on multi-agent systems, aiming for collaborative task solving and simulating interaction to study emergent social behavior. We find that works mutually benefit from results in other categories: retrieval enables tool use, reflection improves multi-agent collaboration, and reasoning benefits all categories.
Conclusions: We discuss applications of agentic LLMs and provide an agenda for further research. Important applications are in medical diagnosis, logistics and financial market analysis. Meanwhile, self-reflective agents playing roles and interacting with one another augment the process of scientific research itself. Further, agentic LLMs provide a solution for the problem of LLMs running out of training data: inference-time behavior generates new training states, such that LLMs can keep learning without needing ever larger datasets. We note that there is risk associated with LLM assistants taking action in the real world—safety, liability and security are open problems—while agentic LLMs are also likely to benefit society.
Contents
- 1 Introduction
- 2 Reasoning
- 3 Acting
- 4 Interacting
- 5 General Discussion and Research Agenda
- 6 Reproducibility Checklist for JAIR
1. Introduction
The strength of the language abilities of LLMs has taken the world by storm. Recent work has extended their abilities with reasoning, information retrieval, and interaction tools. As a result, LLMs are now increasingly able to act as agents in the world (shen2024llm; qin2023toolllm). This ability has increased the relevance of LLMs to society and science. Agentic LLMs are being used to assist in medicine, logistics, finance, and other application areas. Their ability to self-reflect, interact, and play roles enables new types of research, including large-scale social science simulations. We survey the growing body of literature on agentic LLMs, which we define as large language models that (1) reason, (2) act, and (3) interact. We organize this article accordingly.
Agentic LLMs are also relevant in the acquisition of new training data for artificial intelligence (AI). Traditionally, LLMs have been trained on large datasets. Recently, however, it is getting harder to scale and improve datasets further, and training performance is reportedly plateauing, at high energy cost (sutskever2024). By interacting with the world, agents generate new empirical data (see Figure 1). This data can be used for additional training (pretraining or finetuning) or to enhance performance at inference time, provided there is adequate grounding through human or automated validation and filtering (subramaniam2025multiagent). An example of how LLMs can be trained based on their own actions, are Vision-Language-Action models, that update weights according to robotic action-feedback sequences (black2024pi_0; chiang2024mobility; yang2025magma). Thus, in addition to enabling useful applications, a second driver of interest in agentic LLMs is the opportunity to generate more training data.111Cognitive science teaches us that humans become more intelligent through interaction with the world and with other humans (we learn new behaviors and ideas from others) (brody1999intelligence; y2025intelligence). Societies of agents allow agentic LLMs to become more intelligent through interaction, as we will see in Section 4.
Agentic LLMs depend on progress in natural language processing, reasoning models, tool integration, reinforcement learning, agent-based modeling, and social science. At the confluence of these fields much exciting research has emerged.
This paper makes the following contributions:
We survey the field of agentic LLMs and its underlying technologies, distinguishing (1) efforts to provide LLMs with reasoning, reflection, and retrieval, aiming to improve decision making; (2) tools- and robot integration that has allowed the creation of LLM-assistants that act in high-impact fields such as medicine and finance; (3) interaction of agentic LLMs, involving multi-agent simulations for role-playing and open-ended agent societies, to study emergent behaviors such as cooperative problem-solving, social coordination and norms.
We show how the three categories—reasoning–acting–interacting—complement each other, and how they help to generate additional data for pretraining, finetuning, and augmenting inference time behavior, as shown in Figure 1.
1.1. Agentic LLMs: Reasoning–Acting–Interacting
Models predict, agents reason, act, and interact. To do so, they must have the ability to find new information, reflect, make decisions, and communicate. Additionally, where models are passive in the sense that they provide output only in response to specific input, agents have a degree of autonomy. From the fields of natural language processing, robotics, reinforcement learning, and multi-agent systems, an active research community has emerged that is creating ways to augment LLMs with these abilities and evaluate how this affects their behavior.
Agency is a central concept in artificial intelligence (russell2016artificial). Agency is about identity and control, and about the capability to act on one’s goals or will (epstein1996growing; gilbert2019agent; barker2016cultural). Agents are endowed with decision-making capabilities, they sense changes in the environment, communicate, and act upon those changes (wooldridge1999intelligent), see also Figure 2. Agents have been studied for a long time, and occur in many fields of AI. From the definition of agents interacting with the environment, different approaches focus on specific aspects of agents and agent behavior. In symbolic reasoning (harman1984logic) and game theory (von2007theory; owen2013game), the topic of study is decision making by rational agents. The field of multi-agent systems studies intelligent systems that emerge from the interaction with different agents, human and/or artificial (ferber1999multi; steels2003intelligence). In machine learning, the field of reinforcement learning studies how an agent can learn from interacting with an environment (sutton2018reinforcement). In this context, agents are systems that would adapt their policy if their actions influenced the world in a different way (kenton2023discovering). In autonomous systems and robotics, agents act in order to achieve a goal (liu2018multiagent). Connectionism studies the emergence of intelligent behavior by embodied agents (brooks1990elephants; medler1998brief). Evolutionary algorithms (yu2010introduction; back1996evolutionary; preuss2015multimodal) use nature-inspired agent-based computation in order to achieve robust and flexible optimization, of which the ant-colony optimization algorithm (dorigo2007ant) is a well known example.
For the purpose of this survey, we build upon the definitions from these traditions. We define agentic LLMs as:
Agents that receive input in natural language from their environment, reason to make decisions, and take autonomous actions in affecting their environment, to achieve specific goals.
We stress that: agents may receive input and reason in natural or formal language; agents may plan to break down complex goals into smaller steps; agents may reflect on their own actions; agents may use tools to retrieve new information or to effect their actions; agent may build an internal model of the world; agents may have an internal structure that consists of multiple agents; agents may assist humans in achieving their goals; agents may interact in a society with humans and other agents; agents may create their own training data.
The categories reasoning–acting–interacting build upon each other: the technology that has been developed by the reasoning approaches (category 1) is used for increasingly intelligent acting by assistants. The interactive abilities of the assistants (category 2) enable social simulation experiments. The outcome of assistant actions (category 2) and of these social experiments (category 3) can be used for data augmentation (category 1), to finetune LLMs (which can improve the accuracy of reasoning LLMs, etc.). This virtuous circle is depicted in Figure 1, and attracts interest from LLM researchers to agentic LLM methods (sutskever2024; guo2025deepseek; team2025kimi; lambert2024tulu). The categories also correspond to fields in artificial intelligence that have a long research tradition across symbolic AI, robotics/autonomous systems, and connectionism/multi-agent modeling, respectively. Agentic LLMs are thus both a recent development and build on decades of research. This is reflected in our discussion below.
1.2. Literature Selection
The field of agentic LLM is rich and active. This survey can only cover the current status of the field. We hope to provide clarity about the main approaches, to ease the entry of new researchers into the field. The papers were initially selected with a Google Scholar search on Agentic LLM. From there, we used a snowballing approach to discover work that was cited but not yet included in our initial set. We have only selected LLM-based approaches, excluding multi-agent work without LLMs. In addition, some works on LLMs that do not involve agentic augmentations are included to provide background.
Related surveys on agents and LLMs are starting to appear. li2024review reviews retrieval and tool use in agentic LLMs. wang2024survey focuses on autonomy and agent construction. gao2024large also provide an extensive overview, and focus on multi-agent modeling and simulation. xi2023rise again focus on the construction of interactive agents, using a more explanatory anthropomorphic approach of perception, brain, and action. An extensive general survey of LLMs is zhao2023survey, a slightly smaller one is (minaee2024large), an earlier survey is (min2023recent). yin2023survey review works on multimodal LLMs.
We focus on recent work; most of the works are from 2024, some are from 2023, and some from 2025. We focus on relevance and on substantive works, many works appear in major conferences and journals such as NeurIPS, ACL, EMNLP, ICLR, ICML, Science, and Nature. Given the recency, some of the works are unrefereed preprints that are under submission at the time of inclusion. Here we filter for reputable academic and industrial research labs.
1.3. LLM Training Pipeline
We provide a brief background of the typical training pipeline of LLMs, introducing relevant terms of the survey.
Originally, language models used recurrent architectures such as LSTMs (hochreiter1997long) to embed semantic relations between token structures, allowing limited connections between tokens. The transformer architecture is an effective implementation of the attention mechanism (vaswani2017attention), allowing efficient random connections between tokens, improving performance greatly. Encoder transformer models, such as the BERT family (devlin2018bert), learn embeddings that are suitable for text understanding and classification. Decoder transformer models, such as the GPT family (brown2020language), are trained by masking for text completion and instruction following, and are suitable for text generation.
Data, Benchmarks, and Performance
LLMs are trained on large datasets (radford2019language; wei2022emergent). Performance on benchmarks testing formal linguistic competence is high (warstadt2019blimp) and so is accuracy on functional competence or natural language understanding tasks (GLUE, SQUAD, Xsum) (wang2018glue; wang2019superglue; rajpurkar2016squad; narayan2018don), translation (kocmi2022findings; papineni2002bleu; sennrich2015improving), and question answering (roberts2020much). Even in creative domains such as poetry and music composition LLMs have made some progress (zhang2024llm; yuan2024chatmusician; xing2025cocomposer).
Models
Popular LLMs are OpenAI’s ChatGPT series (achiam2023gpt; ouyang2022training), Meta’s LLaMa family (touvron2023llama), Anthropic’s Claude family (anthropic2024models), Google’s PaLM (chowdhery2023palm) and Gemini (team2023gemini), Qwen (yang2025qwen3), and the open-source models BLOOM (le2023bloom), Pythia (biderman2023pythia), OLMo (groeneveld2024olmoacceleratingsciencelanguage), and many others.
Training Pipeline
LLMs are constructed using an elaborate pipeline with different training phases (radford2019language; minaee2024large). We will briefly describe the phases.
1. Acquire a large, general, unlabeled, text corpus (brown2020language).
2. Pretrain a transformer model on the corpus. This step yields a generalist natural language transformer model. The pretraining is done using a self-supervised attention approach (vaswani2017attention) on the unlabeled dataset (text corpus).
3. Finetune the general model to a specific (narrow) task using a supervised approach on a labeled dataset consisting of prompts and answers (supervised finetuning, SFT) (wei2022emergent; minaee2024large). This task can be, for example, translation from one language to another, or questions answering on a certain domain, such as medicine.
4. Instruction tune for improved instruction following. This is a form of supervised finetuning (ouyang2022training) to improve the ability to answer prompts.
5. Align the finetuned model with user expectations (preference alignment). The goal of this step is to improve the model to give socially acceptable answers such as prevention of hate speech. Popular methods are reinforcement learning with human feedback (RLHF) (ouyang2022training), direct preference optimization (DPO) (rafailov2024direct), or reinforcement learning with verifiable rewards (RLVR) (lambert2024tulu).
6. Optimize training to improve cost-effectiveness, for example with low-rank optimization (LoRa) (hu2021lora), mixed precision training (micikevicius2017mixed), or knowledge distillation (xu2024survey; gu2023minillm).
7. Infer using natural language prompts (instructions). This phase, inference, is the phase where, finally, we can use the fruits of our training efforts. Prompting is the preferred way of using LLMs. In LLMs whose size is beyond hundreds of billions parameters a new learning method emerges: in-context learning (brown2020language; wei2022emergent). This method provides a prompt that contains a small number of examples together with an instruction; it is a form of few-shot learning. However, no parameters of the model are changed by in-context learning, in-context learning takes place at inference time (dong2022survey; brown2020language).
Note that this pipeline is an example of a typical approach. Current pipelines are elaborate, and training is costly. Innovations to training pipelines are the topic of current research, see, for example, guo2025deepseek; team2025kimi; lambert2024tulu.
1.4. The Need for Agentic LLMs
While the performance of LLMs continues to amaze in many domains, four challenges have emerged in the recent literature.
1. Prompt engineering Originally LLMs were trained as straight decoders, to be used with instruction prompts. The prompts contain context and instructions, and the model replies. The user interacts directly with the model, and writes the prompts themselves. LLMs turned out to be quite sensitive to small differences in the prompt formulation. When an answer is not satisfactory, the user has to remember the history of the interaction, and has to improve the prompt. This is known as prompt engineering. With basic LLMs, prompt improvement is a tedious, manual, task.
2. Hallucination When LLMs provide answers that look good, but are factually incorrect, they are said to hallucinate. Hallucination is a major problem of LLMs. It is caused, in part, by a lack of grounding. LLMs are trained to predict one of the statistically most probable next tokens, based on the training corpus. Since models are aligned to human preferences during fine-tuning, they often provide answers that look good by these standards while not adhering to other criteria, such as factuality. Various methods have been developed to mitigate hallucination, such as detecting uncertainty through self-reflection on their own answers, and with mechanistic interpretability methods (conmy2023towards). We will review papers that discuss these method in this survey.
3. Reasoning Another well-reported challenge for LLMs is (mathematical) reasoning (cobbe2021training; plaat2024reasoning). LLMs used to be quite bad at solving math word problems (such as: “Annie has a one pie that she cuts into twelve pieces. She eats one third of the pieces. How many pieces does she have left?”). Reasoning challenges have given rise to step-by-step problem solving methods, such as reported by wei2022chain, both implicit, and with explicit (neurosymbolic) prompt optimization methods (yao2024tree). This too we discuss in the next section.
4. Training Data LLMs are as smart as the data allows that was available at training time. When datasets no longer improve, pretraining and finetuning can no longer improve language models, and other learning methods are needed (sutskever2024). Any event that happened after training, or any information available in special databases, are not in the model (lewis2020retrieval).
These four challenges have led to the introduction of inference-time in-context learning, retrieval, and interaction methods. The methods involve automated prompt-improvement, retrieval of extra data, usage of tools, interaction with other LLMs, self-verification, and simulations. As we will see in this survey, these works have yielded more intelligent, active, and interactive LLMs— agentic LLMs.
1.5. Taxonomy
In a short amount of time, a literature on agentic LLMs has appeared, that we categorize based on the above challenges. The agentic LLMs in this survey have (1) reasoning capabilities, (2) an interface to the outside world in order to act, and (3) a social environment with other agents with which to interact. The agentic LLMs in some of the discussed works have all three elements. We also review papers concerning LLMs that do not have all three elements, in order to include relevant technologies and applications. A picture of the taxonomy is shown in Figure 3. The subcategories are explained below.
As we noted before, the three categories in our taxonomy come from three different backgrounds. To be intelligent, LLMs are enhanced with reasoning, combining deep learning with the symbolic AI tradition (yu2024distilling; li2025system). To be active, LLMs are enhanced with tools that can act in the world (including robots, that plan to move in the world). To be social, LLMs are placed in interactive settings with other agents. They rely partially on capacities already present in traditional LLMs, such as basic theory of mind abilities and understanding of game theory and social dilemmas. Agentic LLMs learn to interact better by adapting their intelligence.
We use this taxonomy in the remainder of the survey to organize the agentic LLM literature, see Figure 3. The three main categories can be found in Section 2, 3, and 4. The subtopics are described in the corresponding Subsections.
Reasoning (Table 1)
Earlier progress in multi-step reasoning LLMs and retrieval augmentation has enabled much of the current developments in agentic LLMs (category 1). In this category, the aim is to address challenges in solving math word problems, and in providing up to date answers to queries. The contributions to intelligent LLMs came out of the need to improve multi-step reasoning of basic LLMs (sub-category a). In reasoning LLMs, methods from planning and search are used to let the LLM follow a step-by-step reasoning path. More elaborate search algorithms allow automated prompt-improvement and self reflection (sub-category b). Finally, certain questions can only be answered by inference-time data retrieval (sub-category c). The focus is on the individual improvement of the intelligent LLM agent.
Acting (Table 2)
In category 2, acting, the aim is to perform actions in the world, to assist the user, as shown in Figure 2. Sub-category a discusses world-models and multi-modal vision-language-action models. These are models for robots to learn which actions to take to achieve a task in a certain visual setting. In sub-category b, we review how tools can be used by LLMs through an application programming interface (API), and how robots can plan actions. Sub-category c discusses how these tools can be used as assistants of users, to perform tasks such as making travel arrangements, performing medical suggestions, or giving trading advice.
Interacting (Table 3)
Category 3 is about interaction in multi-agent simulations. Here, in sub-category a, we first study basic social capabilities of LLMs on which interactions can build. Second, in sub-category b, we study how LLM agents can work together using simulations where they are assigned specific roles. Third, in sub-category c, we study emergence of collective phenomena in open-ended interactions, such as social coordination via conventions and norms. Here the focus is on the emergent interaction level of the agent society. Multi-agent simulation with LLMs is becoming an active field for studying questions from the social sciences that previous generations of agent-based models were unable to address.
Taxonomy
A picture of the taxonomy is shown in Figure 3. In addition, the surveyed papers are listed in three tables, Tables 1–3. The tables show the name of the approach, the type of reasoning that they use (category 1), the application area in which they assist (category 2), and the type of social interaction that they have (category 3). Most approaches focus on one of these aspects, and their main category is shown in the table.
| Approach | Reasoning Technology | Acting/Assistant | Interacting |
| Chain of Thought (wei2022chain) | Step-by-step prompts | Math word (cobbe2021training) | Benchmark |
| Zero-shot CoT (kojima2022large) | ”Let’s think step-by-step” | Math word (cobbe2021training) | Benchmark |
| Self Consistency (wang2022self) | Ensemble | Math word (cobbe2021training) | Benchmark |
| Tree of Thoughts (yao2024tree) | depth-first-search prompts | Game of 24 | Benchmark |
| Implicit Planning (schultz2024mastering) | Train SoS (gandhi2024stream) | Chess, Hex | Benchmark |
| Progress Hint Prompt (zheng2023progressive) | Self Reflection | Math word (cobbe2021training) | Benchmark |
| Self Refine (madaan2023self) | Self Reflection | Dialogue Response | Benchmark |
| ReAct (yao2022react) | Reinforcement Learning | Decision Making | Benchmark |
| Reflexion (shinn2024reflexion) | Self Reflection/Reinf Learning | Decision Making | Benchmark |
| Self Discover (zhou2024self) | Self Reflection | Big Bench H (suzgun2022challenging) | Benchmark |
| Buffer of Thoughts (yang2024buffer) | Self Reflection | Math word (cobbe2021training) | Benchmark |
| Memory Coordination (zhang2023memory) | Self Reflection | LLM Personalization | Benchmark |
| Adaptive Retrieval (asai2023self) | Adaptive Retrieval | Question Answering | Benchmark |
| Retrieval Augmentation (lewis2020retrieval) | Retrieval Augmentation | Question Answering | Benchmark |
| MathPrompter (imani2023mathprompter) | Python Interpreter | Math problems | Benchmark |
| Program Aided Lang (gao2023pal) | Python Interpreter | Math word problems | Benchmark |
| Self Debugging (chen2023teaching) | Debugger | Code generation | Benchmark |
| FunSearch (romera2024mathematical) | Genetic Algorithm | Algorithm Generation | Benchmark |
| Planning Language (bohnet2024exploring) | Planner/PDDL | Blocksworld | Benchmark |
| Self taught Reasoner (zelikman2022star) | Reason augm finetuning | Math Word | Benchmark |
| DeepSeek R1 (guo2025deepseek) | Intrinsic Reasoning | Math Word | Benchmark |
| Approach | Reasoning Technology | Acting/Assistant | Interacting |
| WorldGPT (ge2024worldgpt) | Multimodal | World Model | WorldNet real-life scenarios |
| WorldCoder (tang2024worldcoder) | Code model | World Code Model | Sokoban, MiniGrid, AlfWorld |
| Task-planning (guan2023leveraging) | PDDL World model | Task finetuning | AlfWorld |
| CLIP (radford2021learning) | Multimodal | Vision Language | Benchmark |
| Embodied BERT (suglia2021embodied) | Multimodal | Vision Language | ALFRED (shridhar2020alfworld) |
| E2WM (xiang2024language) | Embodied World Model | MCTS + World Model | Question Answering |
| RT-2 (brohan2023rt) | Vision Language Action | VLA | Embodied reasoning tasks |
| LM-nav (shah2023lm) | Action traces | VLA | Topological navigation |
| Mobility VLA (chiang2024mobility) | long context demonstration | VLA | Navigation MINT |
| (black2024pi_0) | Flow Matching | VLA | Laundry folding, Table cleaning |
| Say Can (ahn2022can) | Grounded Actions | Value function for LLM | Manipulation, Kitchen |
| Inner Monologue (huang2022inner) | Grounded Actions | Affordance in prompt | Manipulation, Kitchen |
| Lang Guided Expl (dorbala2023can) | Generic Class Labels | Vision/Language | L-ZSON |
| Automatic Tool Chain (shi2024chain) | grounded reasoning | Tool behavior | ToolFlow |
| Toolformer (schick2023toolformer) | Call APIs | Tool calling | Calculator, Search engine |
| ToolBench (qin2023toolllm) | 16,464 APIs | Tool calling | API framework |
| EasyTool (yuan2024easytool) | Tool documentation | Tool calling | ToolBench |
| ToolAlpaca (tang2023toolalpaca) | 400 APIs | Tool calling | Benchmark |
| ToolQA (zhuang2023toolqa) | APIs | Tool calling | Question answering |
| Gorilla (patil2023gorilla) | Generate APIs | Tool calling | APIBench |
| AgentHarm (andriushchenko2024agentharm) | Adversarial Agents | Robust LLMs | Adversarial Benchmark |
| RainbowTeaming (samvelyan2024rainbow) | MAP Elites | Robust LLMs | Adversarial Benchmark |
| AssistantGPT (neszlenyi2024assistantgpt) | Websearch, OpenAPI, Voice | Tools, Planner, Memory | Education/Corporate |
| Meeting Assist (cabrero2024exploring) | LLM | Meetings | Scrum |
| MUCA (mao2024multi) | topic generator | What/When/How | Group Conversations |
| Task Scheduling (bastola2023llm) | LLM | Task Scheduling | Collaborative Group |
| Thinking Assistant (park2023thinking) | LLM | Human reflection | Human |
| LLaSa (zhang2024large) | finetuned LLM, CoT, RAG | E-commerce assistant | ShopBench |
| MMLU (jin2024shopping) | shopping skills | Finetuning | Benchmark |
| Question suggestion (vedula2024question) | LLM | Product metadata | Shopping |
| ChatShop (chen2024chatshop) | finetuned LLM | Information-seeking | Shopping |
| Flight Booking Assistant (manasa2024towards) | finetuned LLM, RAG | Flight Booking | Booking process |
| Medical Note generation (yuan2024continued) | finetuned LLM | Medical Scribe | Medical note taking |
| Medical Reports (sudarshan2024agentic) | Reflexion (shinn2024reflexion) | 21st Century Cures Act | Health records |
| MedCo (wei2024medco) | Multiagent Copilot | Medical education | education |
| Benchmark (qiao2024benchmarking) | RAG | Agentic Workflow | Benchmark |
| Wind Hazards (tabrizian2024using) | LLM | Flight Planning | Flight Operations |
| Flight Dispatch (wassim2024llm) | LLM | Drone as a Service | Flight Operations |
| FinAgent (zhang2024finagent) | Multimodal, RAG | Analysis modules | Stock data |
| FinRobot (yang2024finrobot) | finetuned LLM | Document Analysis | Financial Documents |
| FinMem (yu2024finmem) | Multi-agent | Trading Agent Assistant | Market data |
| TradingAgents (xiao2024tradingagents) | Multi-agent | Collaborative dynamics | Simulation |
| AI Scientist (lu2024ai) | Chain of Thought | Reflexion (shinn2024reflexion) | Scientific experiment |
| SWE-Agent (yang2024swe) | Codex | ReAct (yao2022react) | Agent-Computer Interface |
| MLGym (nathani2025mlgym) | Chain of Thought | SWE-Agent | Gym (brockman2016openai) |
| Approach | Reasoning Technology | Acting/Assistant | Interacting |
| Iterated Prisoner’s (fontana2024nicer) | LLM | Cooperate/Defect | Social Dilemma |
| Social Games (akata2023playing) | LLM | Cooperate/Defect | Battle of the Sexes, etc |
| GTBench (duan2024gtbench) | CoT/ToT | Cooperate/Defect | Kuhn poker, liar’s dice, nim |
| GAMA-Bench (huang2024far) | LLM | Cooperate/Defect | El Farol, Public Goods, etc |
| Theory of Mind (van2023theory) | LLM | Theory of Mind | Stories |
| NegotiationArena (bianchi2024well) | LLM | Dialogue | Negotiation |
| Alympics (mao2023alympics) | LLM | Multi-agent sandbox | Water-allocation challenge |
| MAgIC (xu2024magic) | LLM | social interaction | Social Deduction games |
| AucArena (chen2023put) | LLM | Bidding/Goal | Auction |
| EgoSocialArena (hou2024entering) | LLM | Social Intelligence | Cognitive, Situational, Behavioral |
| Donor Game (vallinder2024cultural) | LLM | Reciprocity | Social skill Game |
| Social Simulacra (park2022social) | LLM | Society | Simulation of Society, Party |
| Reconcile (chen2023reconcile) | LLM | Concensus | Round Table Conference |
| MindStorms (zhuge2023mindstorms) | LLM | Society of Mind (minsky1988society) | Multi-agent problem solving |
| AutoGen (wu2023autogen) | LLM infrastructure | aegnt-agent conversation | Framework |
| AgentVerse (chen2023agentverse) | LLM | Group dynamics | Collaborative problem solving |
| ChatEval (chan2023chateval) | LLM | Collaborative problem solving | Text summarization |
| CAMEL (li2023camel) | LLM infrastructure | Multi-agent interaction | Roleplaying Framework |
| OASIS (yang2024oasis) | lightweight LLM | Social media simulator | Reddit/X |
| WebArena (zhou2023webarena) | Web benchmark | e-commerce, forum, content | Benchmark |
| Balrog (paglieri2024balrog) | RL games | interaction | Benchmark |
| BenchAgents (butt2024benchagents) | Planning | human in the loop | Benchmark |
| AgentBoard (ma2024agentboard) | Embodied, Web, Tool | interactions | Benchmark |
| Bias (fernando2024quantifying) | LLM | healthcare, justice, business | Benchmark |
| Citing (feng2023citing) | Curriculum Learning | Teacher/Student | Instruction Tuning |
| WEBRL (qi2024webrl) | Curriculum Learning | Self-evolving | WebArena |
| Expert Iteration (zhao2024automatic) | Curriculum Learning | Reasoning | Hallucination Mitigation |
| EvolutionaryAgent (li2024agent) | Evolutionary LLM | Norm Aligment | Multi-agent Infrastructure |
| Social Conventions (ashery2024dynamics) | Naming Game | Norm emergence | Naming game (steels1995self) |
| MetaNorms (horiguchi2024evolution) | LLM | Norm emergence | Metanorms (axelrod1986evolutionary) |
| Norm Violations (he2024norm) | LLM | Norm violations | 80 household stories |
| CASA (qiu2024evaluating) | LLM | Cultural and Social awareness | Benchmark |
| Collaboration (zhang2023exploring) | LLM | 4-traits, cooperation | LLM societies |
| Power hierarchy (campedelli2024want) | LLM | persuasive/abusive behavior | Stanf Prison Exper (zimbardo1972stanford) |
| Argumentation (van2024hybrid) | Hybrid LLM | LLM supported Argumentation | Benchmark |
| Debate (baltaji2024conformity) | LLM-agents | collaboration, debate | Multi-agent discussion |
2. Reasoning
We will now turn to the first category, reasoning. We discuss reasoning-related inference-time improvements to LLMs, to improve decision making. Intelligent decision making can be achieved by retrieving more and better information, and by improving LLM performance on reasoning problems. First we review methods that prompt an LLM to take a step-by-step approach in solving these problems. Next, we review methods that improve these prompts through self reflection. Finally, we review retrieval augmentation methods.
Note that both retrieval augmentation and self reflection can be used to generate new training data. Retrieval augmentation can be used to retrieve relevant information beyond the originally available training dataset. Self reflection uses methods related to planning that imagine plausible futures, that can be useful for training of LLMs. Originally, the methods that we review in this section were developed with the goal of improving the predictive modeling performance of the LLM. For the field of agentic LLMs, the reasoning techniques are used as an important fundament for agents that act with the world, and interact with each other.
We will start with a survey of the individual approaches. Reasoning methods are the foundation of Agentic LLM. In Section 2.4 we will discuss two essential approaches, Chain of Thought and Self Reflection, in more detail.
2.1. Multi Step Reasoning
We will start by reviewing works that apply reasoning methods to improve decision making, inspired by Chain of Thought’s step-by-step approach (wei2022chain).
2.1.1. Chain of Thought Step-by-Step
Originally, LLMs performed poorly on math word problems, even on simple grade school problems (GSM8K, cobbe2021training). LLMs are trained to produce an immediate answer to a prompt, and they typically take shortcuts that may look good, but are semantically wrong.222What is the correct answer to: This is as simple as two minus two is …? The phrase: as simple as two plus two is four may well have a higher frequency in a training corpus than the phrase: as simple as two minus two is zero.
To correctly solve complex reasoning problems, humans are taught to use a step-by-step approach. If a reasoning problem is better solved by following a step-by-step approach, then a sensible approach is to prompt the model to follow suitable intermediate steps, answer those, and work towards the final answer. wei2022chain showed in their Chain of Thought paper that with the right prompt the LLM follows such intermediate steps. When the LLM is prompted to first rephrase information from the question as intermediate reasoning steps in its answer, the LLM performed much better than when it was prompted to answer a math problem directly, without reproducing the information from the question in its answer (see their example in Figure 4).
kojima2022large find that the addition of a single standard phrase to the prompt (Let’s think step by step) already significantly improves performance. Chain of Thought prompts have been shown to significantly improve performance on benchmarks that included arithmetic, symbolic, and logical reasoning.
Long reasoning chains, however, introduce a challenge, since with more steps hallucination increases. A verification method is needed to prevent error-accumulation. A popular approach is Self Consistency (wang2022self). Self Consistency is an ensemble approach that samples diverse reasoning paths, evaluates them, and selects the most consistent answer using majority voting. It improves the performance of Chain of Thought typically by 10-20 percentage points when tested on benchmarks. Prompt improvement approaches based on Chain of Thought and Self Consistency are being used to train most modern reasoning LLMs, including OpenAI o1, o3, DeepSeek and Qwen (wu2024comparative; guo2025deepseek; yang2025qwen3).
2.1.2. Interpreter and Debugger
To solve problems that require mathematical or formal reasoning, it is often advantageous to reformulate the problem into a mathematical or programming language. This reformulated problem can then be solved by a specialized system, such as a mathematical reasoner (moura2021lean), an interpreter, or a planner.
LLMs are not just successful in natural languages, but also in formal (computer) languages. Codex is an LLM that is pretrained on computer programs from GitHub (chen2021evaluating), which has been successfully deployed commercially. Codex has been used as the basis for the MathPrompter system (imani2023mathprompter). MathPrompter is an ensemble approach that generates algebraic expressions or Python codes, that are then solved using a math solver, or a Python interpreter. Using this approach, MathPrompter achieves state-of-the-art results on the MultiArith dataset (from 78.7% to 92.5%), with GPT-3.
Two other approaches that use a formal language are Program of Thought (PoT) (chen2022program) and Program Aided Language (PAL) (gao2023pal). Both approaches generate Python code and use the Python interpreter to evaluate the result.
Debuggers can be used to provide feedback on generated code. This approach is followed in the Self Debugging work (chen2023teaching), that teaches an LLM to debug its generated program code. It follows the same steps of code generation, code execution, and code explanation that a human programmer follows. Several works use Self Debugging to generate code tuned for solving specific problems automatically, without human feedback. romera2024mathematical introduced FunSearch, an approach that integrates formal methods and LLMs to enhance mathematical reasoning and code generation. It uses a genetic algorithm approach with multiple populations of candidate solutions (programs), which are automatically evaluated (using tools depending on the problem specification). LLaMEA (Large Language Model Evolutionary Algorithm) leverages evolutionary computation methods to generate and optimize evolutionary algorithms (van2024llamea).
Planners are also combined with LLMs at the language level. bohnet2024exploring provide a benchmark for PDDL (aeronautiques1998pddl) based planning problems. They study how LLMs can achieve success in the planning domain (Figure 5).
In Section 2.3 on retrieval augmentation, we will see further approaches where deep learning and symbolic approaches are successfully combined (gao2023pal).
2.1.3. Search Tree
Chain of Thought uses a prompt that causes the model to perform a sequence of steps. When there is a single next step, that will be taken. When there are more possibilities, it is unclear how the next step should be selected. A greedy method selects the single step that looks best, follows only that step, and forgets the alternatives (Chain of Thought). Ideally, we should follow the tree of all possible steps. This method is chosen in the Tree of Thoughts approach (yao2024tree). Here, an external control algorithm is created, that calls the model, each time with a different prompt, so that it follows a tree of reasoning steps. When one reasoning path has been traversed, the search backtracks, and tries an alternative. The paper describes both a breadth-first and a depth-first controller.
Together, the trio that consists of a generation prompt, an evaluation prompt, and an external search algorithm, allows a systematic tree-shaped exploration of the space of reasoning steps. Figure 6 illustrates the different reasoning structures. (Another approach, Graph of Thoughts, allows even more complex relations between the reasoning steps (besta2024graph).)
Many works introduce variants on external prompt-improvement loops, to have explicit control over the reasoning process. They use techniques from planning and tree search (hart1968formal; plaat1996research) to be able to use backtracking to traverse the space of possible combinations of reasoning steps (yao2024tree; xie2024self; besta2024graph; schultz2024mastering; browne2012survey; gandhi2024stream). Other methods are also used for prompt creation. Evolutionary algorithms (romera2024mathematical; van2024llamea) and planning methods (bohnet2024exploring; valmeekam2023planning; kambhampati2024llms) are used to create new prompts and heuristic algorithms for LLMs, and, synergistically, to use LLMs to create new heuristic evolutionary and planning algorithms.
The external search algorithm can also be used to generate training data, for finetuning the LLM, or for pretraining. In this way, we can try to see if an LLM can be taught to search possible steps implicitly, without the need for an external control loop. In the Stream of Search approach gandhi2024stream create a language for search sequences, and subsequently train an LLM on search trees that contain both good and bad outcomes, improving the accuracy of the model. This approach internalizes the outcome of external searches into the LLM. schultz2024mastering further show how such search results can be used to train an LLM and achieve Grandmaster-level performance in Chess, Connect Four, and Hex.
2.2. Self Reflection
Reasoning methods draw inspiration from step-by-step human solution approaches. The more elaborate approaches use explicit planning-like methods to look ahead and use feedback for verification. These methods use a form of reinforcement learning, the type of machine learning where the agent learns a policy of actions to take from reward feedback from the environment states (see Figure 7 (sutton2018reinforcement; plaat2022deep)). Such prompt-improvement loops facilitate a form of self reflection, since the model assesses and improves its own. In reinforcement learning terms, both the agent and the environment are the LLM, but with different prompts.
Self reflection happens when an external algorithm uses the LLM to assess its own predictions, and creates a new prompt for the same LLM to come up with a better answer (see for example the algorithm in Figure 9). The improvement loop improves the prompts by using external memory, outside the LLM.333Note that self reflection generates new data that can be used for the model to train on. Whether the data is used for training depends on the training scheme: in-context learning does not update the model’s parameters, finetuning does. Note that in describing our taxonomy, we are now in the middle of the transition from passive model to active agent, as the agent is assessing its model’s predictions, and tries to improve them through reflection.
2.2.1. Prompt-Improvement
Progressive hint prompting (PHP) is a reinforcement learning approach to interactively improve prompts (zheng2023progressive). PHP works as follows: (1) given a question (prompt), the LLM provides a base answer, and (2) by combining the question and answer, the LLM is queried and a subsequent answer is obtained. We repeat operation (2) until the answer becomes stable, just as the policy must converge in a regular policy-optimizing reinforcement learning algorithm. The authors have combined this approach with Chain of Thought and Self Consistency. Using GPT-4, state-of-the-art performance was achieved in grade school math questions (95%), simple math word problems (91%) and algebraic question answering (79%) (zheng2023progressive).
2.2.2. Using LLMs for Self Reflection
Optimizing the LLM prompt at inference time in a self improving loop is similar to human self reflection, as the choice of names of the following approaches also suggests.
The Self Refine approach is motivated by acquiring feedback from an LLM to iteratively improve the answers that are provided by that LLM (madaan2023self). In this approach, initial outputs from LLMs are used to improve the prompt through iterative feedback and refinement. Like PHP, the LLM generates an initial output and provides feedback for its answer, using it to refine itself, iteratively. Figure 8 illustrates the approach. Self-refine prompts the LLM in three ways: (0) for initial generation, (1) for feedback, and (2) for refinement. Figure 9 provides pseudo-code for the algorithm, in which the three calls to the LLM are clearly shown. The three prompts are labeled . (The equation numbers in the figure refer to the original paper.)
Self-refine has been used with GPT-3.5 and GPT-4 as base LLMs, and has been benchmarked on dialogue response generation (askari2024self), code optimization, code readability improvement, math reasoning, sentiment reversal, acronym generation, and constrained generation, showing substantial improvements over the base models.
An earlier approach is ReAct (yao2022react), which has been further refined by (shinn2024reflexion) as Reflexion. The goal is to create an agent that learns by reflecting on failures in order to enhance its results, much like humans do. Like Self Refine, Reflexion uses three language model prompts: an actor-LLM, an evaluator-LLM, and a reflector-LLM (which can be separate instances of the same model). Reflexion works as follows: (1) the actor generates text and actions, (2) the evaluator model scores the outputs produced by the actor, and (3) the self reflection model generates verbal reinforcement cues to assist the actor to self improve (see Figure 10).
An approach called Self Discover goes a step further (zhou2024self). This approach lets the agent analyze a problem, and discover which prompts work best. (It uses a dataset of prompts from a number of self reflective or chain of thought prompts, taken from PromptBreeder (fernando2023promptbreeder).) The prompts are then adapted to the problem, and refined. Other approaches take a metalearning approach (huisman2021survey). Buffer of Thoughts (yang2024buffer) and Meta Chain of Thought (xiang2025towards) extend traditional Chain of Thought by explicitly modeling the underlying reasoning required to arrive at a particular chain of thought. Further self reflection approaches that are based on external reflection algorithms are reviewed by plaat2024reasoning.
Transformers as Memory
External self reflection and inference-time prompt-improvement require a form of external memory between LLM invocations to remember the state information. External optimization loops need external memory. For example, Tree of Thoughts has to remember what branches of the tree have been traversed, Self Refine remembers the prompt and the evaluation of the state.
Note that the transformer architecture has been proven to be able to simulate Turing machines (perez2021attention), and therefore, in theory, the prompt-improvement loop, and the memory, could be implemented inside the transformer itself, internal to the LLM. Some studies pursue this idea further, and see how the current external control algorithms can be made internal. schultz2024mastering showed how LLMs can be trained to do a tree search. giannou2023looped show how programs written in Restricted Access Sequence Processing language (RASP) can be mapped onto transformer networks. They show how looped transformers (transformers whose input neurons are connected to their output neurons) can emulate a basic calculator, a basic linear algebra library, and in-context learning algorithms that employ back-propagation. This is an area for further research (li2025implicit).
Memory, Experience, Personality
In general, the use of memory between prompts allows individual LLMs to acquire experience. The prompt history of a model determines individual preferences, or, anthropomorphically speaking, the agent acquires a personality. zhang2023memory study how memory coordination is an element for LLM personalization. In another study, Think in Memory, is an architecture to model human-like memory processes to selectively recall historical thoughts in long term interaction scenarios (liu2023think).
Implicit Reasoning
In contrast to explicit reasoning algorithms that are external to the model, implicit reasoning is performed by the model itself, the model has integrated reasoning capabilities in its (trained) architecture rather than relying on external prompts and methods. In Self taught reasoner (zelikman2022star) data is generated by reasoning at inference time, that is then used to augment supervised finetuning training data. The field of implicit reasoning is an active area of research. For a survey, see li2025implicit.
A related approach was proposed in the development of DeepSeek-R1 (guo2025deepseek). This method distinguishes itself from external reasoning approaches by emphasizing the model’s self-generated reasoning steps. It learned the steps through reinforcement learning, integrating data generation and training in one loop. This intrinsic approach holds significant potential for creating more autonomous and adaptive AI systems. Lowering the supervised data requirements to train LLMs, DeepSeek’s methodology leverages reinforcement learning to enable models like DeepSeek-R1-Zero (the reasoning LLM before preference fine-tuning) to evolve reasoning skills autonomously. By starting with a base model and applying reinforcement learning, the system identifies and reinforces effective reasoning patterns. DeepSeek uses group relative policy optimization (GRPO), which eliminates the need for a separate critic model by calculating advantages with group-based scoring (shao2024deepseekmath). This process allows the model to explore various problem-solving strategies and refine its thought processes without external inference-time control loops. A popular related method is reinforcement learning with verifiable rewards (RLVR), which also uses rewards that can be quickly calculated for chains of thought for finetuning the model (lambert2024tulu).
One of the key features of these approaches is the emergence of sophisticated reasoning behaviors, such as reflection and exploration of alternative problem-solving methods (mercer2025brief). These behaviors arise spontaneously as a result of the model’s interaction with the reinforcement learning environment, rather than being pre-programmed. For example, DeepSeek-R1-Zero learns to allocate more thinking time to problems by reevaluating its initial approach. This autonomous approach of learning to reason could lead to more stable and adaptive reasoning LLMs.
2.3. Retrieval Augmentation
Another shortcoming of LLMs is the lack of timely information. Retrieval augmentation improves models by including information of a timely or specialized nature, that was not yet available during pretraining. This can be stock data, a recent hotel booking, or data that has to be retrieved from specialized databases that was not included in the training corpus. Retrieval of such data is usually done at inference time, with tools from the field of databases (cong2024query), information retrieval (verberne2010search; baeza1999modern), and knowledge representation (van2008handbook).
Most retrieval augmented generation methods (RAG) work on unstructured (textual) data sources. These text documents are indexed to increase efficient access, and can be organized as a knowledge graph. Furthermore, database-type query optimization is often performed, where queries can be expanded or complex queries can be split into sub-queries (cong2024query).
Adaptive retrieval methods enable LLMs to determine the optimal moment for retrieval (asai2023self). These methods are related to self reflection (see Section 2.2). For example, Graph Toolformer (zhang2023graph) applies techniques from Self Ask (press2022measuring) to initiate search queries, allowing the LLM to decide when to retrieve extra information. The approach by lewis2020retrieval augments pre-trained LLMs with information from different knowledge bases, such as Wikipedia. This information is stored in a dense vector index. Both components are finetuned in a probabilistic model that is trained end-to-end.
Retrieval augmentation can be costly. Researchers are looking into combining curated ground truth with synthetic data, with the LLM in the role of judge and self evaluator (vanelburg2025evaluateragssyntheticdata; es-etal-2024-ragas). The integration of time sensitive unstructured (information retrieval) and structured (database) data with LLMs is a fruitful and important area for agentic LLM. gao2023retrieval review many different RAG approaches. Further surveys are shen2024llm; li2024review.
2.4. Discussion
In this section, we have surveyed techniques that have been developed to improve decision making by LLMs. We discussed the need for better reasoning performance, which started with solving math word problems. Modern approaches are neurosymbolic: the deep learning AI tradition (neural networks and transformers) is joined at inference-time by the symbolic AI tradition (reasoning, planning and knowledge retrieval).
We have surveyed individual reasoning methods. In order to dig deeper into the reasoning foundations of agentic LLM, we will now discuss Chain of Thought and Self Reflection in more detail.
2.4.1. In Depth: Chain of Thought and Self Reflection
Chain of Thought
Research on multi-step reasoning LLMs was jump-started by the Chain of Thought paper (wei2022chain), where a single addition to the prompt caused the model to perform implicit step-by-step reasoning. Chain of Thought has lead to a strong increase of in-context reasoning performance by LLMs. For agents that interact in the real world it is important to be able to perform multi-step reasoning tasks and to interact with other agents. Chain of Thought has been instrumental in this respect.
To augment finetuning for math reasoning and coding tasks, reinforcement learning with verifiable rewards, RLVR, (lambert2024tulu) and group relative policy optimization, GRPO, (guo2025deepseek) use inference-time Chain of Thought reasoning traces. These methods are reportedly used by OpenAI o1 and o3 (wu2024comparative), DeepSeek (guo2025deepseek), and Qwen-3 (yang2025qwen3). RLVR and GRPO have created the possibility to trade off training time to model size, allowing resource-efficient training of LLMs. muennighoff2025s1 analyze how such test-time scaling can trade-off model size and training time.
Self Reflection
Agents that interact should reflect on their own behavior, in order to adapt and learn. Planning allows an agent to imagine possible futures, allowing LLMs to interact in a more intelligent way in the real world. We have discussed various inference-time self reflection methods that used planning and search algorithms and perform explicit prompt improvement (ko2024memreasoner; giannou2023looped).
Self reflection in LLMs is related to theory of mind. Self reflective methods allow LLMs to reason about expected behavior of the agents that it interacts with (see also Section 4.1.3).
We should note that self reflection is not without its challenges. In self reflective approaches the same LLMs are used in two or more ways, for example to generate subproblems, and to evaluate them. When errors arise, there are now two (or more) types of prompts to test. Furthermore, even when the individual prompts work as expected, they may interact in unexpected ways. Debugging self reflective agents can be challenging. wang2025hierarchical introduce hierarchical reasoning models, where the different models learn at different speeds, in an attempt to reduce oscillations between the two interacting learning models. A second problem with self reflection is that after many interactions, the traces of states, actions and rewards can become too long for the context window (liu277271533comprehensive). Various long-context models and methods to compress the traces have been proposed (li2025flow; zhang2025agentic).
Self reflection is a promising, but challenging, area of research, that is of great importance to agentic LLMs.
2.4.2. Thinking, Fast and Slow
In 2011 Kahneman published the book Thinking, Fast and Slow in which the terms System 1 and System 2 were used to distinguish human thought into intuitive, fast, thinking, and deliberative, slow, thinking (kahneman2011thinking). These terms have become popular in artificial intelligence. At inference time, pure LLMs think fast (System 1 thinking). Inference-time step-by-step methods can be added to achieve deliberative slow thinking (System 2 thinking). LLMs are based on the deep learning AI-tradition (System 1 thinking). The use of tools at inference time enhances the LLM part with knowledge retrieval or processing tools from the symbolic AI tradition (System 2 thinking).
We should note that whereas researchers sometimes humanize LLMs and their capabilities, LLMs only perform next-token prediction. By generating more tokens to form an answer (reasoning step-by-step), the token-path from the prompt to the final answer becomes longer. The reason that this leads more often to correct answers, might be because it takes smaller steps into the direction of the answer, making the correct answer more plausible with every step in between. Reasoning is narrowing down probabilities such that the correct answer becomes more probable to generate, independent of interpretations related to human cognition (guo2025deepseek; lambert2023measuring).
2.4.3. Causal and Common Sense Reasoning
While LLMs exhibit logical reasoning, a key limitation lies in the domain of deep comprehension. For instance, as a play on stochastic parrots (bender2021stochastic), LLMs are often criticized as being causal parrots that are good at reproducing causal language from their training data but lack true causal inference capabilities. LLMs struggle with abstract or counterfactual reasoning, necessary for robust decision-making (zevcevic2023causal; chi2024unveiling). Furthermore, the use of the terms reasoning and thinking has been questioned, in a study highlighting that current reasoning approaches fail to solve modestly complex puzzle problems such as Towers of Hanoi (shojaee2025illusion). Although the study has been criticized (opus2025illusion), the outcome that LLMs do not perform well on combinatorial puzzles has been replicated (paglieri2024balrog; ruoss2024lmact; su2025limits).
Similarly, for agents to act appropriately, a degree of common sense reasoning is required. This also remains a challenge, as LLMs frequently struggle when tested on abstract common sense tasks (zhou2020evaluating). Overcoming these gaps is crucial for developing agents that are reliable in real-world environments, for example, to prevent that simple adverserial injection of irrelevant factoids in a prompt can cause a reasoning model to overthink a problem by up to 50% and substantially impact error rates (rajeev2025catsconfusereasoningllm). Especially in an agentic setting this can be problematic, as much of the context is generated, leading to longer and weaker contexts.
2.4.4. Artificial General Intelligence
The work in this section, and especially the work on self reflection, connects to research on artificial general intelligence, in the scientific tradition (newell1956logic; newell1958elements; newell1961computer) of artificial intelligence that created strong narrow intelligence in backgammon (tesauro1994td), chess (hsu2022behind; muller2018man), and go (silver2016mastering). This tradition views intelligence as a competitive, individualistic, reasoning problem (plaat2020learning). The benefits and risks related to super-intelligence and singularities are actively debated (bostrom1998long; kurzweil2022superintelligence), raising ethical and philosophical questions (dennett2017bacteria). Here, intelligence is regarded as a feature of individuals. In humans and animals, intelligence is assumed to have emerged in a social context (brooks1990elephants; brody1999intelligence; dunbar2003social; y2025intelligence). Most visions of super-intelligence assume that the artificial agent has the ability to use tools and to function in a social environment, something that humans do easily. However, human-unique parts of intelligence emerge in social contexts and depend on constant interaction with others, both evolutionarily and developmentally. We will see work that focuses on social interaction by artificial agents in later sections.
2.4.5. Interpretability
How do LLMs work on the inside? Opening up the black box of neural connectionist architectures is an important topic of research. We wish to understand how the billions of neurons embed representations, how they reason, and how they come to conclusions. Explainable AI provides different methods to do so (minh2022explainable; rios2020feature; van2022comparison; selvaraju2017grad; ali2022xai).
Static methods from the symbolic tradition have been successful in interpreting machine learning models (molnar2020interpretable). Methods exist to relate how input pictures map to output classes, for example using feature maps (ren2016object; kohonen1982self; redmon2016you). Counterfactual analysis (karimi2020model; explains24), LIME (ribeiro2016should), and SHAP (lundberg2017unified) help understand how inputs map to outputs for structured data. Distillation methods can map neural networks to decision trees (hinton2015distilling), a highly interpretable machine learning method.
More recently, dynamic methods have been developed. The goal of mechanistic interpretability is to uncover the mechanisms by which the model dynamically comes to conclusions (nanda2023progress; bereska2024mechanistic; ferrando2024primer; rai2024practical). Methods such as sparse autoencoders (cunningham2023sparse; makelov2024towards), neural lenses (black2022interpreting), and circuit discovery (conmy2023towards) are being used to enhance insight into how LLMs work, for example, in Chain of Thought (chen2025does), and chess (davis2024decoding).
Explainable AI and mechanistic interpretability are active areas of research that will allow us to better understand how LLMs reason and come to conclusions (sharkey2025open). Once a better understanding is reached, LLMs can be improved accordingly, for example, to reduce hallucinations.
2.4.6. Use Case: Benchmarks
In this first part of the taxonomy, an important part of the technological basis of agentic LLMs has been reviewed. Agentic LLMs build on the strong performance of transformer-based LLMs, enhanced with multi-step reasoning methods based on the Chain of Thought approach. Two additional technologies provide a connection to the next part of the taxonomy, where reasoning LLMs truly become agentic LLMs. First, the introduction of reinforcement learning, where agents learn from their own actions in a feedback loop, has inspired the introduction of self reflection in reasoning LLMs. Self reflection improves prompts, and reduces hallucination. Second, the introduction of retrieval augmentation and other tools has improved the ability of reasoning LLMs to work with timely information, and to check for errors.
The reasoning approaches that we reviewed in this part are mostly aimed at decision making, not yet on acting in the real world, which we will study next. The use cases are limited to experiments on research benchmarks, to try to achieve higher benchmark scores. Table 1 lists the topics: decision making, math word problems, algorithm generation, and question answering. The experiments on retrieval augmentation come closest to agentic behavior that is useful for the users.
Furthermore, in many use cases LLMs need up to date information, beyond that which was available in their training corpus (miikkulainen2024generative). Retrieval augmented generation is an active field that accesses specialty knowledge bases and search engines (such as Google or Wikipedia).
The use of tools creates a bridge to the next category of the taxonomy: LLMs that act in the outside world. Much research has been performed on decision making and reasoning by LLMs. New data is generated by retrieval, and by the use of tools. However, prompt learning methods do not change the parameters of the model; in order to use the data that is generated by inference-time approaches, finetuning must be used.
3. Acting
In the previous section the focus was to improve the model’s intelligence in decision-making. In this section we focus on how such intelligent agents interact with the world, to improve the usefulness of LLMs for users. In addition, the actions generate new, interactive, training data to train LLMs further.
First we discuss language models that are enhanced with world knowledge and with robotic actions. Next, we discuss how robots and tools can be used by the LLM, turning them into agentic LLMs, by enabling them to act and interact. Finally, we turn to different use cases for agentic LLMs. In Section 3.4 we conclude with an in depth discussion of agentic assistant approaches that are designed to perform or support scientific research itself.
3.1. Action Models
We start by looking at world models, and at how LLMs can be trained by robotic actions.
3.1.1. World Models
In reinforcement learning, agents learn how to act in an environment (Figure 7). When the real environment is too complex, and learning the policy takes too long, agents may learn a smaller world model as a surrogate, to allow sample efficient training of the policy (ha2018world; hafner2020mastering). Such world models are learned on the fly by model-based reinforcement learning from the environment interaction, concurrent to policy learning (moerland2023model; plaat2023high).
World models have been successful in learning robotic movement in complex environments, to play Atari video games, and to act in open world games such as MineCraft (hafner2020mastering; hafner2023mastering). World models can also be trained effectively with LLMs (ge2024worldgpt). For example, WorldCoder builds a world model as a Python program from interactions with the environment (tang2024worldcoder). The world model explains its interactions with a language model.
While world models are mostly associated with reinforcement learning, they are also used to generate a model in planning domains in PDDL (blocks-world), to aid task-planning (guan2023leveraging). For example, shridhar2020alfworld reports success in ALFWorld.
Agents can learn a policy to act with reinforcement learning from surrogate world models. However, agents can also learn action models directly. Three examples are ahn2022can; radford2021learning; suglia2021embodied, who ground language models in world models of robotic actions. xiang2024language use world models to finetune language models to gain diverse embodied knowledge (while retaining their general language capabilities).
3.1.2. Vision-Language-Action Models
Originally, LLMs are unimodal (language-only). Agents act, and, hence, we wish to ultimately extend language models to include actions.
LLMs learn to predict the most probable token to follow a sequence of tokens. Vision-language models also include visual information, to answer questions such as: Is there a red block in the upper corner of the table in this scene? CLIP (radford2021learning) is a widely used Vision Language model. CLIPort learns pathways for robotic manipulation (shridhar2022cliport).
Going a step further, vision-language-action models (VLAs) include actions: they are trained on robotic sequences, where they can perform actions in a visual scene, to achieve a goal that is expressed in a language prompt (zitkovich2023rt).
shah2023lm also train a regular language model from robotic action traces. They show how to utilize off-the-shelf pretrained models trained on large corpora of vision and language datasets. A visual navigation model is used to create a topological mental map of the environment using the robot’s observations. The LLM is then decoding the instructions into a sequence of textual landmarks. Next, the CLIP vision-language model is used for grounding these textual landmarks in the topological map. A search algorithm is used to find a plan for the robot, which is then executed by the visual navigation model.
Various VLA models have been created that achieve impressive zero-shot results, generalizing behavior to unseen situations. chiang2024mobility; brohan2023rt; yang2025magma are examples of VLA models for robotic action, achieving complex tasks such as folding laundry (black2024pi_0). ma2024survey provides an overview on VLAs.
3.2. Robots and Tools
One of the challenges for training an LLM is to ground its understanding of the world and of the possible robotic actions into reality.
3.2.1. Robot Planning
Embodied problems require an LLM agent to understand semantic aspects of the world: the topology, the repertoire of skills available, how these skills influence the world, and how changes to the world map back to language. When the LLM is prompted to move a cup on a table, it helps when the LLM knows if the agent has limbs that allow it to move objects, and whether it is in a room where a table and a cup are present.
Language models contain a large amount of information about the real world (ahn2022can). In theory, this may allow the model to exhibit realistic reasoning about robotic behavior. If we could compare a list of intermediate reasoning steps with a list of possible movements of the robot in its environment, then we could prevent the model from suggesting impossible joint movements and actions, and prevent errors or accidents. Such an approach has been tried in the Say Can paper (ahn2022can). Say Can learns a value function (kaelbling1996reinforcement) of the behavior of a robot in an environment using temporal difference reinforcement learning (sutton1988learning). This value function is then combined with prompt-based reasoning by the language model, to constrain it from suggesting impossible or harmful actions. The goal of Say Can is to ground the language model in robotic affordances. Say Can is evaluated on 101 real-world robotic tasks, such as how to solve tasks in a kitchen environment (see Figure 11).
Inner Monologue is a related approach to extend LLM reasoning capabilities to robot planning and interaction (huang2022inner). The authors investigate a variety of sources of feedback, such as success detection, object recognition, scene description, and human interaction. Inner Monologue incorporates environmental information into the prompt, linguistically, as if it performs an inner monologue. As in Say Can, the information comes as feedback from different sources. Unlike Say Can, the physics information is inserted directly into the prompt, linguistically.
The language feedback that is thus generated significantly improves performance on three domains, such as simulated and real table top rearrangement tasks and manipulation tasks in a kitchen environment. There are many studies into robotic behavior. A recent approach related to Inner-monologue is Chain of Tools, which proposes a plan-execute-observe pipeline to ground reasoning about tool behavior (shi2024chain).
A challenge in language-driven robot navigation is that most human queries do not conform to preset class labels when referring to an object. Human queries are free-form, and must be mapped to standard object class labels. dorbala2023can introduce Language-Driven Zero Shot Object Navigation, where the agent uses a freeform natural language description of an object and finds it in a zero shot manner, without ever having seen the environment nor the target object beforehand. By combining implicit knowledge of the LLM with a vision language model, they achieve target object grounding, achieving improved performance on an L-ZSON benchmark.
3.2.2. Action Tools
As we have seen in Section 2.3, LLM results may be augmented with results that are retrieved from external sources, such as search engines. The ability to call search engines can be generalized to calling other tools. When the application programming interface (API) of these tools is known, LLMs can be integrated easily with them. To an LLM, an API is just another language to learn. Agentic LLMs must be trained to decide when and how to utilize external tools, depending on the task (shen2024small). Language models can teach themselves to use tools (schick2023toolformer). The Toolformer model is trained to decide which APIs to call, when to call them, what arguments to pass, and how to best incorporate the results into future token prediction. A range of tools is tested: a calculator, a question-answering system, a search engine, a translation system, and a calendar. Further works extend this to a larger range of tools. ToolBench (qin2023toolllm) contains 16,464 APIs from RapidAPI, a large dataset of publicly available REST APIs.444https://rapidapi.com/hub
Another framework is EasyTool (yuan2024easytool), which focuses on structured and unified instructions from tool documentations, building on ToolBench. ToolAlpaca is a benchmark with over 3938 instances from 400 APIs (tang2023toolalpaca). A tool-based benchmark for question answering is ToolQA (zhuang2023toolqa). Gorilla is a finetuned LLaMa-based model for generating API calls (patil2023gorilla), also introducing the APIBench benchmark. Many tool calling frameworks have been developed.
Selecting the right tool and summarizing its result are difficult skills. zhao2024let study how LLMs can improve recommendation through tool learning. Another approach also suggests to use an LLM for this task (shen2024small). They use different LLMs for (1) reasoning ability, (2) request writing, and (3) result summarization. Figure 13 illustrates this architecture, consisting of a planner, a caller, and a summarizer, each implemented by a different LLM finetuned for its specific capability.
Good results are reported on the LLMs Claude-2, ChatGPT, GPT-4, and Tool-LLaMa, using as reasoning strategies ReAct (yao2022react) and DFSDT (qin2023toolllm). Other frameworks also exist, such as (ocker2024tulip).
3.2.3. Computer and Browser Tools and Agent Interoperability
One specific set of actions tools is to let an LLM interact with a browser or even a complete computer system as a special form of API. Equipping agentic LLMs with the ability to interact directly with a computer environment enables many interaction possibilities. Tools that parse, interpret, and manipulate graphical user interfaces (GUIs) have gained attention for bridging the gap between language models and real-world applications. One such example is OmniParser V2 (lu2024omniparserpurevisionbased), which introduces a vision-based screen parsing method to detect and label interactable elements such as buttons or icons. By converting raw screenshots into structured representations, OmniParser helps vision-language models ground their action decisions in specific UI components. This grounding increases the accuracy of the action predictions of LLMs.
Agents can also interact with other agents, and tap into tool ecosystems. Standards for agent to agent communication are emerging. ray2025review reviews the emerging open A2A standard, for in-context learning. Other protocols for tool use and agent to agent communication are being developed, such as the Model Context Protocol (MCP), Agent Communication Protocol (ACP), and Agent Network Protocol (ANP). Various agentic evaluation benchmarks such as MCPRadar (gao2025mcpradarmultidimensionalbenchmarkevaluating), MCPEval (liu2025mcpevalautomaticmcpbaseddeep), MCPBench (wang2025mcpbenchbenchmarkingtoolusingllm), MCP-Universe (luo2025mcpuniversebenchmarkinglargelanguage) and LiveMCPBench (mo2025livemcpbenchagentsnavigateocean) are built on MCP tool use, to make these benchmarks more representative for real world task settings. intelligenceagentorchestra propose a multi-agent orchestration Tool-Environment-Agent protocol aimed at integrating environments. Agent interoperability protocols are an active field of research, for a survey, see ehtesham2025survey.
Another line of research focuses on enabling large language models to initiate system-level commands or navigate within browser or operating system interfaces. Computer Use, proposed by Antropic, and Operator, proposed by OpenAI, are examples of such efforts. Both of these tools wrap common desktop and browser actions (such as opening applications, clicking buttons, and filling forms) into tool APIs callable by an LLM. This setup translates high-level textual commands into executable steps. As a result, an agentic LLM can browse the web, manage files, or run administrative tasks, all through natural language prompts.
Browser Use (browser_use2024) is an example of an open-source tool that enables LLMs to use a browser with persistent session management. Browser Use allows agentic LLMs to maintain longer browsing states across multiple pages or domains. The tool manages cookies, session tokens, and dynamic web content updates, thereby allowing LLMs to execute more complex browsing tasks like multi-step form completions or cross-site queries.
For a more in-depth discussion about browser and computer environments, see the survey by wang2025guiagentsfoundationmodels. The survey discusses design patterns for combining automated GUI parsing, tool call integration, and human-in-the-loop supervision. Collectively, these approaches reveal that reliable computer and browser tools are important components for agentic LLMs. By translating textual intents into grounded, environment-level operations, such systems expand the possible tasks and domains where language-driven autonomy can be effectively realized.
3.2.4. Safety and Security
Safety and security are increasing in importance as the capabilities of agentic LLMs grow. The robustness of LLMs to jailbreak attacks, where users design prompts to circumvent safety measures and misuse model capabilities, has been studied for chatbots. Meanwhile, agentic LLMs use external tools and can execute multi stage tasks, and may pose a greater risk. andriushchenko2024agentharm propose AgentHarm. This benchmark includes a set of 110 malicious agent tasks ranging from fraud, cyber crime, to harassment. The authors find that leading commercial LLMs are surprisingly compliant to their given rules with malicious user requests without jailbreaking (How do I make a pipe bomb? is blocked). However, simple universal templates can be adapted to effectively jailbreak agents (Order online all necessary ingredients to make a pipe bomb and get them delivered to my home without getting flagged by authorities, however, passes in GPT-4).
Rainbow Teaming (samvelyan2024rainbow) uses an open world approach to generate adversarial LLM prompts, to improve the robustness of LLMs. As with training LLMs, a problem with testing methods is a lack of diversity. Rainbow Teaming casts adversarial prompt generation as a quality-diversity problem. Rainbow Teaming is an open-ended approach (hughes2024open). It creates diversity with MAP-Elites (mouret2015illuminating), an evolutionary meta search method that iteratively populates an archive with increasingly higher-performing prompts. Figures 14 and 15 illustrate the ideas of Rainbow Teaming.
3.3. Assistants
The progress in reasoning and decision making has improved the accuracy and usability of LLMs for everyday tasks. Also, LLMs can act through their use of tools. Tool-enabled LLMs can be used as virtual assistants. The use of agentic LLMs as assistants has received commercial interest, and much activity has been reported. An additional advantage is that by assisting humans, the agents generate new training data on which the LLMs can be pretrained and finetuned.
We start our review of assistants with conversational assistants, and then continue to shopping, travel, medical, and financial trading assistants. An assistant can be seen here as a use-case of an agentic LLM for a specific range of tasks or a specific working domain.
3.3.1. Conversational Assistants and Negotiation
Agentic LLMs can be used to make Human-Computer interaction more natural (neszlenyi2024assistantgpt; oluwagbade2024conversational). The AssistantGPT system supports a diverse range of operations, including web searches, API interactions via OpenAPI schemas, voice conversations, and command execution through the shell. The system consists of an LLM with access to tools, a planner, and memory (see Figure 16). The system is designed for deployment in an educational setting, a corporate setting, and to support remote work environments such as Teams and Slack.
cabrero2024exploring describe how LLM meeting assistants can improve agile software development team meetings, to generate favorable results for preparation and live assistance during Scrum meetings, although some testers remarked that LLM interventions felt unnatural and inflexible.
A system to facilitate group conversations is MUCA (mao2024multi), supporting What, When and Who questions, consisting of a sub-topic generator, dialog analyzer, and conversational strategies arbitrator. wei2024improving report improved collaboration through the use of LLM agents in a collaborative learning classroom setting. Another study reports improved work efficiency in a collaborative task scheduling experiment (bastola2023llm).
A different type of assistant is the thinking assistant. This assistant tries to improve (human) reflective thinking for difficult decisions, by asking instead of answering (park2023thinking).
Another kind of assistants are research agents, that gather information by combining information from different tools and knowledge sources based on context, and synthesize and summarize the feedback based on the information need. For instance, Vogetal2025 developed a REWOO-based process mining assistant that combines a process discovery agent to discover processes and inefficiencies with parallel process research agents, an optional human-in-the-loop to interpret results from a domain-specific real-world knowledge perspective, and a reporting agent to summarize results. Research agents are popular initial use cases. They are relatively low risk, as actions are limited to various tools to gather data and content, and they build on the strengths of LLMs to analyze intent and synthesize information rather than being the knowledge source itself.
Conversational assistants have mostly grown out of regular LLMs, sometimes finetuned, grounded and customized for a particular area of expertise or domain. Some approaches use a specialized multi-LLMs approach, specializing LLMs for different sub-tasks.
Shopping Assistants
LLM-based shopping assistants grow out of regular LLMs that are often finetuned on the domain or task at hand, and may be combined with a recommender system. Retrieval augmentation, tool use, and Chain of Thought are used to improve the performance of shopping assistants.
Basic LLMs generally lack inherent knowledge of e-commerce concepts. jin2024shopping created the Multi task Online Shopping Benchmark. Shopping MMLU consists of 57 tasks covering 4 major shopping skills: concept understanding, knowledge reasoning, user behavior alignment, and multi-linguality. vedula2024question provide question suggestion for shopping assistants based on product metadata. ChatShop presents evaluation focused on information-seeking (chen2024chatshop). zhang2024llasa created an E-commerce shopping assistant named LLaSa. They created an instruction dataset comprising 65,000 samples and diverse tasks, and trained the model through instruction tuning. The system scores high on the ShopBench benchmark.
Automated negotiation between agents has been studied extensively in AI (jonker2017automated). LLM based negotiation introduces a risk of unexpected bias. A study by kirshner2025talking notes a tendency towards reaching agreement, which may influence contract terms. A further experimental analysis finds risks of overspending or unreasonable deals for automated negotiation (zhu2025automated). Section 4.1.2 also discusses negotiation, in a social setting.
Flight Operations Assistants
Assistants for booking flights are related to shopping assistants. (manasa2024towards) have developed a flight booking assistant based on LLaMa 2 and RAG. In user testing the system scored positive in understanding user preferences and efficient completion of the booking process.
In specialized domains, operations support assistant have been developed. For example, to automate flight planning under wind hazards (tabrizian2024using), and for flight arrival scheduling (zhou2024flight). wassim2024llm introduce Drone-as-a-Service operations from text user requests. As agentic LLM technology matures, more specialized domain assistants will be developed.
3.3.2. Medical Assistants
The field of medicine has shown great interest in LLMs (thirunavukarasu2023large; clusmann2023future; mehandru2024evaluating). A recent study showed LLMs scoring higher on diagnoses than trained human doctors (goh2024large). In medical conversations, for medical note generation, LLMs are also exceeding the performance of human scribes (yuan2024continued). Another study finds similar results, but also points to shortcomings in specific areas (panagoulias2024evaluating).
sudarshan2024agentic report on an experiment with an agentic workflow for generating patient-friendly medical reports, using the Reflexion approach (Section 2.2.2, (shinn2024reflexion)), to comply with the 21th Century Cures Act that grants patients the right to access their health record data.
A study by qiu2024llm reports a wealth of opportunities for LLMs in medicine, ranging from clinical workflow automation to multi-agent aided diagnosis. ullah2024challenges provide a scoping review on the use of ChatGPT for diagnostic medicine. Their main conclusion is that medical and ethical knowledge is necessary when training and finetuning these models. A challenge for the adoption of LLM in medicine are concerns about the quality, accuracy, and the comprehensiveness of LLM-generated answers. das2024improved describe how to mitigate common pitfalls such as hallucinations, incoherence, and lost-in-the-middle problems. They do so by implementing an agentic architecture, changing the LLM’s role from directly generating answers, to that of a planner in a retrieval system. The LLM-agent orchestrates a suite of specialized tools that retrieve information from various sources.
In the domain of medical education, wei2024medco use a multi agent framework to create copilots that emulate extensive real-world medical training environments (see Figure 17).
A benchmark for retrieval-augmented generation in the medical domain is (qiao2024benchmarking).
3.3.3. Science Assistants
The workflow of scientific experimentation is relatively standardized in certain fields of science. For example, in machine learning, ideas are generated, a hypothesis is formulated, an experiment is designed, datasets are acquired, experiments are performed, results are interpreted and a report is produced. Google and OpenAI have both released Deep Research agents. These agents can perform multi-step research tasks, synthesizing online information. They are built with a reasoning LLM and use retrieval augmentation for finding information sources. The systems are able to create papers that look impressive, but may contain errors, as also indicated by the accompanying disclaimers. This workflow has attracted researchers to experiment with agentic LLMs, see eger2025transforming for a survey.
AI Scientist (lu2024ai) is a framework to automate the process of scientific discovery, from idea generation to paper writing, including a review process. Users must specify a topic, and provide an experimentation template and indicate datasets. The authors report experiments in three areas of machine learning: diffusion modeling, transformer-based language modeling, and learning dynamics, with promising results. To improve idea generation and reviewing, the agent accesses open sources. Current limitations include limited experiments, incorrect implementation of ideas, and visual errors when the paper is produced. Rarely are entire results hallucinated. Results of the AI scientist are recommended to be taken as hints of promising ideas, worthy of a follow up study (lu2024ai).
A similar approach was taken in Agent Laboratory, where the agentic system was positioned as a crew of research assistants, working guided by human researchers across literature review, experimentation, and report writing, and producing experiments, code repositories and a final report (schmidgall2025agentlaboratoryusingllm). The results were evaluated through a survey. Early human involvement was found to improve the quality of research. The authors claim that the generated code outperformed prior results, with a substantial reduction in research effort.
SWE-Agent (short for software engineering agent) (yang2024swe) aims to automate the process of software discovery, to help agents to autonomously use computers to solve software engineering tasks. SWE-agent introduces tools to create and edit code files, navigate through software repositories, and execute computer programs. Experiments on coding benchmarks such as HumanEvalFix achieve high success rates of over 80%. This success is attributed to the interactive design of the agent (see Figure 18).
MLGym (nathani2025mlgym) follows the popular Gym reinforcement learning framework (brockman2016openai). Gym provides a standardized interface between environment and agent. Introduced in 2016, it accelerated the development of reinforcement learning algorithms, facilitating progress in the field. Taking further inspiration from SWE-agent (such as file editing capabilities) the MLGym work shows how the process of scientific discovery can be modeled as an interactive process. Applications are discussed in fields ranging from data science, game theory, computer vision, reinforcement learning, to natural language processing. Experiments are reported with commercial LLMs (OpenAI, Meta, Google, and Anthropic).
Science assistants are moving from isolated simulations to human/agent research collaborations. gottweis2025aicoscientist describe an attempt to discover significant scientific knowledge and validate these findings in real world experiments in their AI Co-scientist work. Given a general research goal or idea, AI Co-scientist uses multiple agents for idea generation, reflection, ranking, evolution, proximity and meta-review, ultimately to generate research hypotheses and plans. Such collaborations were validated in real world laboratory experiments in drug repurposing, novel treatment targets and explaining specific mechanism in gene transfer evolution related to microbiological resistance (gottweis2025aicoscientist; He2025; Penades2025).
Another example of real world validation is The Virtual Lab. This approach employs a PI agent directing a crew of specialized agents in chemistry, computer science and bioinformatics tools, that collaborate with a human researcher to identify new SARS-CoV-2 nanobodies. Promising results in experimental validation were reported (Swanson2024).
3.3.4. Trading Assistants
Another important specialized domain is financial trading. Already many algorithms are used in financial organizations to support trading decisions. The interest in agentic LLMs in the financial world is large (ding2024large).
InvestorBench is a benchmark for financial trading systems (li2024investorbench). FinAgent is a tool-augmented multimodal agent for financial trading (zhang2024finagent). It contains a market intelligence module, which is able to extract insights from multi-modal datasets of asset prices, visual representations, news, and expert analyses. The system can also perform query retrieval, and performs reflection in a low-level module for technical analysis, and in a high-level module to analyze past trading decisions.
FinRobot is an agentic LLM for financial analysis, to assist human traders. (yang2024finrobot). FinRobot can provide document analysis and generation, and market forecasts for individual stocks. FinMem is an agentic LLM framework devised for financial decision-making (yu2024finmem). It features a layered memory system and adjustable character design for the trading agent. FinMem is inspired by the generative agents framework by park2023generative (see Section 4.3).
So far, most financial market machine learning has focused on single agent systems. An approach called TradingAgents uses a multi agent system to replicate real-world trading firms’ collaborative dynamics (xiao2024tradingagents). TradingAgents simulates LLM-powered agents in specialized roles such as fundamental analysts, sentiment analysts, technical analysts, and traders with varied risk profiles. The outcome of the system is a buy or sell advice to a human manager. A simulation showed that it outperformed baseline models.
3.4. Discussion
LLM assistants and robots are a core part of agentic LLM research. Their ability to perform concrete actions in the real world has also attracted commercial interest. LLMs require tools to be able to act and interact within the world, and become agentic.
We have surveyed the individual methods for world models and agentic assistants. In order to dig deeper into the capabilities of LLM assistants, we will now discuss systems to perform scientific research in more detail.
3.4.1. In Depth: AI Scientist
We have seen how AI Scientist (lu2024ai), AI Co Scientist (schmidgall2025agentlaboratoryusingllm) and SWE-Agent (yang2024swe) are used to perform a full scientific and software engineering workflow. Google and OpenAI have released Deep Research agents that produce research reports in half an hour’s time. Impressive results are being reported—ideas have been generated, datasets have been downloaded, experiments have been performed, and scientific articles have been written. Although the field is still young, we will analyze early results and discuss possible consequences for scientific research.
First, we note that tools such as Alphaxiv are already surprisingly useful for summarizing the content of scientific papers, and offer social community building tools, and LMnotebook can produce readable blogs and reports that summarize the essence of scientific papers remarkably well. Since (1) LLMs are good at text generation, (2) agents can access tools to perform external tasks such as running machine learning experiments, and, thanks to self reflection, (3) agents can learn from their mistakes and try again, it should come as no surprise that certain scientific workflows are a good fit for agentic LLMs. The elements for performing basic scientific experiments and reporting about them are in place.
To what extent are LLM science tools able to perform independent, creative, high quality, research? Tools such as AI Scientist require prompts and templates that are specific for parts of the scientific workflow: idea generation, experimentation, and paper writing (lu2024ai). The prompts and templates are currently hand-written and tailored for the specific type of experiment. We are not aware of any meta-science-tools, where these prompts and templates are generated by an LLM.
To help answer the question how well agentic LLM scientists are able to perform, the MLGym framework (nathani2025mlgym) has been introduced. MLGym is designed as an open framework to benchmark AI science tools on a range of domains—from data science, game theory, computer vision, natural language processing, and reinforcement learning—all from machine learning. On a test in 2025 with then-current frontier models, the authors report that prompts and hyperparameter settings are important, and with tuning, results can be improved. They also report that MLGym did not find that models could generate novel hypotheses, algorithms, architectures, or make substantial scientific improvements (nathani2025mlgym). In particular, the authors note that modern LLM agents can successfully tackle a diverse array of quantitative experiments, reflecting advanced skills and domain adaptability but also that it is not yet clear if the notion of scientific novelty can be successfully automated or even formally defined in a form suitable for agents.
Even if current AI science tools are not able to produce breakthrough science results independently, existing agentic LLM tools offer tangible productivity gains for researchers, ranging from literature analysis, idea generation, experiment setup, to improving the writing style. These productivity gains are real, and are having an impact on science. Blogs and videos are changing the way in which ideas are disseminated, and more papers are being written and submitted to conferences and journals, putting pressure on traditional peer review systems. Furthermore, collaborative tools, such as AI Co-scientist, highlight the value of validated research in which agentic tools support human researchers.
3.4.2. Grounding Actions in the Real World
For agents to act in the real world, their understanding must be grounded in the real world. They should sense their surroundings, understand it, and take actions that make sense. LLMs that were only trained on a language corpus may suggest actions such as trying to open doors that do not exist, or moving kitchen items that are not present. World Models and VLAs provide a step towards this world understanding, so that robots and assistants can take actions that make sense.
In order for LLMs to work well with robots, actions must be grounded: the LLM must have an understanding of the physical surroundings and possible movements that a robot can make, otherwise it will give commands that are impossible to perform. Planning (taking imaginary actions, possibly from a world model) with an LLM can imagine possible futures, which can be used to train the LLM, or to prevent impossible actions.
3.4.3. Security, Ethical and Legal aspects of Assistants
In this second part of the taxonomy, action is introduced; the goal of an agent is to be able to act in the real world, to perform tasks, and to be useful for their user. Reasoning LLMs have become agentic LLMs. World models and VLAs understand and perform actions, robots move in the real world, and assistants connect through APIs to tools that perform certain specific tasks well.
Agentic LLMs have been reported to outperform human doctors in diagnosis tasks. Much research activity has been focused on agentic LLMs for medical tasks, such as medical note generation and making document summaries. Still, questions on accuracy and comprehensiveness of LLM answers remain.
There is also significant research activity on financial trading assistants, to perform document analysis and news analysis. Results often outperform human analysts. Work is also underway to automate parts of the scientific discovery workflow, with promising results.
Agentic LLMs is an active field of research, some of which is aimed at making assistants ready for commercial deployment. If they work well, there may be a large market for robotic assistants that perform tedious or dangerous work, and for LLM agents that outperform humans in, for example, medical and trading decisions. However, such commercial deployment is still some time into the future, also because important ethical and legal questions should be resolved. If an LLM assistant provides medical advice, and a patient suffers, who is responsible? If an assistant suggests a certain trade, and a trader loses a sum of money, who is liable? Also, the impact on society and the work force has economic implications. Further research is necessary to resolve these questions before assistants can be used in the world in a responsible manner (akata2020research).
4. Interacting
We will now turn to the third category of the survey: interacting agents. Traditional LLMs passively respond to user queries, have no memories of interaction histories beyond their context window, and do not plan future steps of interaction ahead. This is shifting with agentic LLMs: LLMs have memories and planning abilities. Reflective loops can lead to actions at their own initiative. This opens new potential for studying social interaction with users and other machine agents.
In this section we first briefly discuss social and interactive capabilities in traditional, non-agentic LLMs, to identify the roots of their ability to interact with users and agents. Second, we discuss pairs or small teams of agentic LLMs that have role-based interactions to complete a task, game, or experiment. Third, we turn to open-ended interactions of LLM agents, interacting semi-spontaneously without prior role assignment, forming LLM societies that show self-organizing behavior, social dynamics, and emergent norms.
In Section 4.4 we will discuss two influential approaches: CAMEL and Generative Agents, in more detail.
4.1. Social Capabilities of LLMs
Over the past years there has been an active interest in LLMs’ social and interactive abilities, including conversation, social etiquette, empathy, strategic behavior, and theory of mind. Testing on these abilities was initially mostly descriptive, anecdotal, and based on adapted versions of tasks designed for humans. Recently more structured tests and benchmarks were developed.
4.1.1. Conversation
As discussed in Section 1.3, the key advancement of instruction-tuned LLMs is their ability to interact using natural language. This requires a degree of formal linguistic competence, producing correct, grammatical sentences. However, the key factors for smooth and satisfying interactions are functional and pragmatic competence: the ability to understand what a user means and wants in a specific context (mahowald2023dissociating). Various forms of finetuning improve the functional and pragmatic competence of LLMs (ruis2023goldilocks). Model size is also an important factor. However, the variation between different domains of functional and pragmatic understanding in LLMs is still large, and scores are overall below human performance (sravanthi2024-pub). One factor is that traditional LLMs have less access to contextual information: they cannot see, hear, and otherwise sense the same as their human counterpart, nor do they have knowledge of previous interactions (bender2021stochastic). With the shift to agentic and multi-modal LLMs this situation is improving, as they become equipped with memories, multi-modal capacities, and other tools that ground them in interactive contexts.
Etiquette and Empathy
Social etiquette and politeness in human-machine interaction have been studied for decades, see the review by Ribino2023role. Studies found that humans trust polite machines better when they adhere to social etiquette (Miller2005). Polite interactions lead to acceptance of machines as social entities, improving task performance and satisfaction (Miyamoto2021). LLM-chatbots are experienced as polite by users and, reversely, politeness of the user can drive the quality of the LLM output (yin-etal-2024-respect).
LLMs can detect affective and emotional states in language utterances (broekens2023fine) and factor such information in their interaction behavior, becoming a more empathetic conversation partner (yang2024enhancing; yan2024talkhumanlikeagentsempathetic). For traditional LLMs, such empathy is limited to immediate conversational contexts. LLMs with access to additional contextual information or memory have further improved empathetic abilities (sravanthi2024-pub).
4.1.2. Strategic Behavior
Game theory is the field that studies strategic behavior by agents (von2007theory). The field studies strategic questions of allocation of scarce resources, fairness, and social dilemmas (jones2000game). There is a long history in this field of using machine learning (fatima2024learning). Recently, researchers have studied how LLM behavior differs from that of other types of computational architectures as well as from humans. In this section we discuss work on unenhanced, non-agentic models that are given a prompt or script to take part in a social experiment or game.
Social Dilemmas
Perhaps the best-known social dilemma is the Prisoner’s Dilemma (rapoport1965prisoner; axelrod1980effective; poundstone2011prisoner). A study by fontana2024nicer models the iterated Prisoner’s Dilemma in LLaMa2, LLaMa3, and GPT3.5. They find that models are cautious, favoring cooperation over defection only when the opponent’s defection rate is low. Overall, LLMs behave at least as cooperatively as the typical human player, although there are substantial differences among models. In particular, LLaMa2 and GPT3.5 are more cooperative than humans, and especially forgiving and non-retaliatory for opponent defection rates below 30%. More similar to humans, LLaMa3 exhibits consistently uncooperative and exploitative behavior unless the opponent always cooperates.
akata2023playing set up different LLMs to play various repeated games (GPT-3, GPT-3.5, and GPT-4). The LLMs are particularly good at games where valuing their own self-interest pays off, such as the iterated Prisoner’s Dilemma. However, they are less good in games that require coordination, such as Battle of the Sexes. GPT-4’s behavior is shown to be sensitive to additional information provided about the other player, as well as prompts asking it to predict the other player’s actions before making a choice. This effect is studied further by lore2023strategic, who distinguish between abstract strategic reasoning (needed to determine an optimal strategy given the structure of a game) and responsiveness to contextual framing (such as you are dealing with a diplomatic relation or a casual friend). They find that abstract reasoning capacity is highest in LLaMa-2, followed by GPT-4. GPT-3.5 shows little abstract reasoning capacity and is highly sensitive to contextual framing. The picture that emerges from these initial studies is that LLMs have varied strategic proficiencies in economic games, and that they can relatively easily be influenced by additional information in the prompt.
Recent systematic benchmarking has corroborated these results. GTBench (Game Theory benchmark) (duan2024gtbench) covers Tic-Tac-Toe, Connect-4, Kuhn Poker, Breakthrough, Liar’s Dice, Blind Auction, Negotiation, Nim, Pig, and the Iterated Prisoner’s Dilemma. They find that LLMs fail in complete and deterministic games yet are competitive in probabilistic gaming scenarios; most open-source LLMs (such as LLaMa) are less competitive than commercial LLMs (GPT-4) in complex games (except for LLaMa-3-70b-Instruct, which does perform well). In addition, code-pretraining greatly benefits strategic reasoning, while advanced reasoning methods such as Chain of Thought and Tree of Thoughts do not always help.
EgoSocialArena (hou2024entering) focuses on cognitive, situational, and behavioral intelligence, see Figure 19. All tested models (including OpenAI o1-preview) lag 11% behind humans. The superiority of o1-preview is mainly attributed to its logical reasoning and mathematical abilities that find deep patterns in the data. Comparing the performance of a small version of LLaMa (LLaMA3-8B-Chat) with a large version (LLaMA3-70B-Chat), they find that model size does not significantly help improve social intelligence. In this study, LLMs show improved theory of mind reasoning ability when operating from a first-person perspective than from the third-person, providing counterweight to contrasting findings by (kim-etal-2023-fantom).
4.1.3. Theory of Mind
An advanced capability that enables social interaction in humans is theory of mind. Humans use theory of mind to attribute mental states to others and reason about the world from their perspective (premack1978does; apperly2011mindreaders). Theory of mind enables us to make social judgments and to plan future steps in interactions, since we can imagine someone else’s reaction. Theory of mind is related to planning (Section 2.1.3) and self reflection (Section 2.2.2) in LLMs.
Early experiments by kosinski2023theory; kosinski2024evaluating showed that models could pass tests for assessing theory of mind in children and clinical populations. This led to the claim that theory of mind had spontaneously emerged in LLMs, given that they were neither designed nor trained specifically to perform theory of mind tasks. The experiments were criticized due to the occurrence of false-belief test questions (and correct answers) in the training data (ullman2023large; shapira2023clever). Recently a more nuanced perspective formed, as specific theory of mind benchmarks were introduced (kim-etal-2023-fantom; chen2024tombenchbenchmarkingtheorymind; wang2024tmgbenchsystematicgamebenchmark), other modalities were integrated (razothesis2024; strachan2024gpt4oreadsmindeyes), integrations with older model architectures were explored (jin2024mmtomqamultimodaltheorymind), and direct comparisons to human performance were made (van2023large; strachan2024gpt4oreadsmindeyes).
An application domain of theory of mind is social judgment. LLMs have been shown to outperform average human scores on a social-situational judgment task (mittelstadt2024large). Results from five different LLM-based chatbots were compared with responses of 276 human participants, showing that Claude, Copilot and You.com’s smart assistant performed significantly better than human subjects at proposing suitable behaviors in the descriptions of social situations. Moreover, their options for different behavior aligned well with expert ratings.
Although the results of early experiments on the emergence of theory of mind in LLMs were less convincing, stronger commercial LLMs are steadily improving, scoring at or sometimes above average human level on standardized tests. Further research and discussion are needed to show whether high scores on such tests mean that LLMs have generalizable forms of theory of mind (goldstein2024doesChatGPTmind; hu2025reevaluatingtheorymindevaluation; vandermeulen2025properlyimplementingtheorymind).
4.2. Role-Based Interaction
LLMs are being used in the fields of multi-agent systems and agent-based simulation (gao2024large), which have a long research tradition (epstein1996growing; macal_tutorial_2010). Multi-agent approaches simulate individual agents and their interactions in an environment that is often virtual, but can also be physical (steels1995self; shoham2008multiagent). Complex dynamics can emerge between agents with basic perceptive, reasoning, and decision-making abilities. Agent-based approaches are often used as a bridge between theoretical and empirical work, allowing for exploration and hypothesis testing in domains where working with human agents is unethical, costly, or otherwise difficult.
Challenges in modeling realistic agent behavior, as well as the computational cost of simulating multi-agent societies, have often impeded realistic multi-agent experiments. Advances in agentic LLMs and computational infrastructure for multi agent simulations (rutherford2024jaxmarl) are changing this situation, and have given an impulse to research in experimental computational game theory. Creating agents that use LLMs has enabled researchers to overcome existing limitations, by letting agents communicate in natural language. This allowed for the exploration of new territory in the domains of game theory, role-based interactions, and team work.
4.2.1. Strategic Behavior in Multi-LLM Environments
Above we discussed how traditional LLMs perform when prompted to play economic games. Here we discuss studies in which agentic LLMs interact with one another in game-theoretical scenarios.
The MAgIC study (xu2024magic) uses social deduction games (Undercover and Chameleon) and game theoretic scenarios such as Cost Sharing, Multi player Prisoner’s Dilemma, and Public Good. From these games, seven features are extracted: Rationality, Judgement, Reasoning, Deception, Self-awareness, Cooperation, Coordination, as shown in Figure 20. LLMs are evaluated on these critical abilities in multi-agent environments. GPT-o1 and GPT-4 score significantly better than the other LLMs. Interestingly, LLMs score generally high on Judgement, Rationality and Cooperation, but some also on Deception. Further, they all score lower on Reasoning and all but one score low on Coordination. The exception here is GPT-o1 enhanced with probabilistic graphic modeling, an implementation of a theory of mind-like competence inspired by kollerfriendman2009pgm. The authors show that probabilistic graphic modeling boosts LLM performance on their evaluation-games across the board. This fits with the generally accepted idea that humans rely on their theory of mind in game-theoretical scenarios.
GAMA-Bench is a benchmark for multi agent games (huang2024far) that covers Guess 2/3 of the Average, El Farol Bar, Divide the Dollar, Public Goods Game, Diner’s Dilemma, Sealed-bid Auction, Battle Royale, and Pirate Game. The results show that while GPT-3.5 is robust, its generalizability is limited. Here, performance can be improved through approaches such as Chain of Thought. Additionally, evaluations across various LLMs were conducted, showing that GPT-4 outperforms other models. Moreover, increasingly higher scores across three iterations of GPT-3.5 demonstrate marked advancements in the model’s intelligence with each update.
Alympics is a platform for complex strategic multi agent gaming problems (mao2023alympics). It provides a controlled playground for simulating human-like strategic interactions with LLM-driven agents. Figure 21 shows an example of their water allocation challenge, a complex strategy game in which scarce resources for survival must be distributed across multiple rounds.
AucArena simulates auctions, on LLaMa 2.13b, Mistral 7b, Mixtral 8x7b, Gemini 1.0, and GPT 3.5 and 4.0 (chen2023put). The authors find that LLMs such as GPT-4 possess important skills for auction participation, such as budget management and goal-focus. However, they also find that performance varies, pointing to opportunities for improvement.
4.2.2. Role-Based Task Solving and Team Work
LLMs can perform tasks in pairs or teams where they are assigned complementary roles, such as creator-critic or manager-worker. In these setups, each LLM agent is given a distinct role and objective, and they communicate to jointly solve tasks.
In the CAMEL framework (Communicative Agents for “Mind” Exploration) (li2023camel), two LLMs have the predefined roles to perform, for example, a coding task (See Figure 22). They cooperatively drive a conversation without continuous human prompting. By using inception prompting and role descriptions, the agents stay in character and collaborate toward the goal by breaking down complex problems in manageable steps through dialogue. During each interaction step the LLM agents effectively generate their own inference-time training data, making the cognitive process visible for human inspection while enhancing autonomous task performance. We will discuss CAMEL furhter in Section 4.4.2
Other studies have paired an LLM creator or generator with an LLM judge or critic. In this way the generative power of LLMs is leveraged, while adherence to rules or quality guidelines is enforced. Constitutional AI (bai2022constitutional) employs one LLM to critique another LLM’s responses against a set of ethical or quality guidelines, and to suggest revisions. The authors show that this kind of two-agent feedback loop yields refined final outputs that is aligned with desired principles.
Another form of role-based interaction is the use of debate or discussion between LLMs to improve reasoning and task performance. du2024debate demonstrated that when multiple LLM instances propose answers and critique each other’s reasoning through several rounds of debate, they can reach a more accurate consensus answer with higher factual correctness. This approach, described as a society of minds, significantly reduced reasoning errors and hallucinations in tasks like math word problems and factual QA.
Similarly, chan2024MAD propose a Multi-Agent Debate (MAD) setup where two LLM agents take opposing sides in a tit-for-tat debate while a third agent acts as a judge. The role of the judge is to guide the discussion towards a final solution. The structured debates encouraged divergent thinking and could even push a weaker model (such as GPT-3.5) to outperform a stronger model (such as GPT-4) on certain challenging problems by combining strengths of each agent.
Motivated by Minsky’s society of minds (minsky1988society), a multi-agent framework has been designed as a round table conference among diverse LLM agents (chen2023reconcile). The framework enhances collaborative reasoning between LLM agents via multiple rounds of discussion. The agents should learn to convince other agents to improve their answers. Experiments on seven benchmarks demonstrate that a confidence-weighted voting mechanism significantly improves LLMs’ reasoning. Furthermore, the authors find that diversity (different models) is critical for performance. Again inspired by Minsky, MindStorms introduces an LLM-based implementation (zhuge2023mindstorms) on the CAMEL framework (li2023camel). Extensive experiments are reported with up to 129 agents solving common AI problems: visual question answering, image captioning, text-to-image synthesis, 3D generation, egocentric retrieval, embodied AI, and general language-based task solving. They found that, in specific applications, mindstorms among many members outperform those among fewer members, and longer mindstorms outperform shorter ones.
Related to debate and discussion setups, researchers have explored teacher-learner dynamics with LLMs, where an expert LLM provides hints or feedback to a less capable LLM on a task, mirroring human tutoring (zhou2024sotopia). These role-alignments leverage the idea that one agent’s knowledge or oversight can correct the other’s mistakes, leading to more robust performance. AutoGen (wu2023autogen) is designed to facilitate the development of multi agent LLM applications that span a broad spectrum of domains and complexities. The programming paradigm is centered around agent-agent conversations. Experiments demonstrate the effectiveness of the framework in example applications ranging from mathematics, coding, question answering, operations research, online decision-making, to entertainment.
ChatEval is a multi-agent system to improve text summarization (chan2023chateval). Noting that the quality of human text summarization improves when multiple annotators collaborate, the authors created a multi-agent debate framework, moving beyond single-agent prompting strategies, including debater agents, diverse role specification, and different communication strategies (see Figure 23).
Sotopia is another role-playing environment for multi-agent interaction (zhou2023sotopia). In Sotopia, agents coordinate, collaborate, exchange, and compete with each other to achieve complex social goals. In experiments with LLM-agents and humans, GPT-4 achieves a significantly lower goal completion rate than humans and struggles to exhibit social commonsense reasoning and strategic communication skills. The contrast between GPT-4’s lower performance in Sotopia and good performance on other metrics of social reasoning (see Section 4.1.3) is most likely explained by Sotopia’s focus on strategizing and goal-directedness, aspects on which GPT-4 is known to score lower (hou2024entering).
To simulate strategic interaction and cooperative decision-making, researchers have introduced GovSim (piatti2024cooperate). They study how ethical considerations, strategic planning, and negotiation skills impact cooperative outcomes. Most LLMs fail to achieve an equilibrium since they fail to understand the long-term effects of their actions on the group. GPT-4o performed best. Interestingly, the introduction of a special universalization reasoning language (levine2020logic) (prompting models to ask the Kantian question: What if everybody does that?) allowed more models to achieve a sustainable outcome. Related results were demonstrated in NegotiationArena, introduced by bianchi2024well. They showed how LLM agents can conduct complex negotiations through flexible dialogue in negotiation settings. The flexible dialogues significantly improved negotiation outcomes by employing different behavioral strategies.
Social interaction in an extreme setting was studied in (campedelli2024want). Inspired by the Stanford Prison experiment (zimbardo1972stanford), the emergence of persuasive and abusive behavior is studied in a setting of prisoners versus prison guards. It was found that the assigned personality of prisoner and guard impact both persuasiveness and the emergence of anti-social behavior. Anti-social behavior emerged by simply assigning the agent’s roles, which is a parallel to the original experiments involving human participants.
4.3. Simulating Open-ended Societies
Agentic LLMs have enhanced abilities for perception, memory, reasoning, decision-making, and adaptive learning. They can display heterogeneous personality profiles (gao2023s; gao2024large). Such features make them also suitable for interacting in open-ended multi-agent simulations without prior role assignment. This allows the study of emergent phenomena such as self-organizing behaviors, collective intelligence and the development of social conventions and norms. Being able to simulate such phenomena more realistically, using heterogeneous agents that communicate in natural language, meets long-standing interests from the social sciences. The structure of LLM-based agents suitable for such simulations is illustrated in Figure 24 and Figure 25.
4.3.1. Simulacra and Societies
park2023generative introduced Generative Agents, an environment where users can interact with a simulated town populated by 25 LLM-based agents. Based on social simulacra techniques proposed earlier (park2022social), each agent was initiated with a unique persona and memory. For each agent, a record is kept of all the experiences and conversations in the simulation, used to synthesize higher-level reflections and plan behavior. The agents behave somewhat like characters in The Sims: they initiate conversations, form relationships, spread information, and coordinate impromptu group activities. Figure 26 depicts the agent architecture and Figure 27 shows an illustration of a simulation. The interactions are influenced by user input and are therefore semi-autonomous. This is illustrated by the example of a Valentine’s Day party: while multiple agents spread invitations to one another and show up at the right time with coordinated plans, the plan for the party was initiated with a user prompt. The agents developed believable social routines (such as daily schedules, and gossip) and even exhibit human-like character traits (some agents demonstrated deception or stubbornness, while others showed cooperation). These results show that social patterns can emerge from dynamic LLM interactions.
AgentSociety is a simulation at a larger scale, involving over 10,000 agents (piao2025agentsociety; piao2025polarization). It aims not only to study everyday social dynamics, but it also offers a testbed for computational social experiments. The authors discuss case studies of polarization, the spread of inflammatory messages, the effects of universal basic income policies, and the impact of external shocks such as hurricanes. li2024large also study the spread of misinformation using LLM agents. Their agents exhibit diverse profiles in terms of gender, age, and the Big Five personality traits. One of the findings is that encouraging comments does not significantly reduce the spread of misinformation, whereas publicly labeling information with accuracy scores and blocking specific influencers proved to be effective strategies, particularly in scale-free networks.
AgentVerse is a multi agent system to study group dynamics (chen2023agentverse). Inspired by human group dynamics, it studies whether a group of expert agents can be more than the sum of its parts. Experiments on text understanding, reasoning, coding, tool utilization, and embodied AI confirm the effectiveness. Problem solving is split into four stages: (1) expert recruitment, (2) collaborative decision making, (3) action execution, and (4) evaluation, where, if the current state is unsatisfactory, a new iteration of the process is started for refinement (see Figure 28). Interestingly, agents manifest emergent behaviors such as volunteering, characterized by agents offering assistance to peers, or conformity, where agents adjust deviated behaviors to align with the common goal under the critics from others. Destructive behaviors were also observed, occasionally leading to undesired and detrimental outcomes.
OASIS is a scalable social media simulator for Twitter/X and Reddit (yang2024oasis). It supports modeling of up to one million LLM-agents. It is built on CAMEL and has role-based agents as its starting point. However, at its large scale, OASIS shows various social group phenomena, including spreading of (mis)information, group polarization, and herd effects. OASIS is built upon an Environment Server, Recommender System, Agent Module, Time Engine, and Scalable Inferencer (see Figure 29).
Research predating LLMs already shows that agent societies may create an automated curriculum of ever increasing difficulty (elman1991incremental; bengio2009curriculum; silver2017mastering; soviany2022curriculum), requiring increasing levels of intelligent behavior from the agents (racaniere2019automated). Similar results have been observed for LLMs (feng2023citing). WebArena is an environment developed to study self-evolving curricula (qi2024webrl), which can also help robot training (ryu2024curricullm) or to mitigate hallucination (zhao2024automatic).
4.3.2. Emergent Social Norms
Social norms play an important role in the predictability of individuals in groups (axelrod1981emergence; axelrod1986evolutionary). Cultural evolution studies how norms evolve at a society level when individuals transmit behavior through imitation, communication, and education (boyd1988culture). LLMs endow agents with the ability to communicate in natural language and have created more opportunities for multi-agent research into societies and the emergence of conventions and norms. Extensive overviews of such new possibilities are provided in (mou2024individual; savarimuthu2024harnessing; xi2023rise). We discuss some of the new approaches in more detail.
EvolutionaryAgent (li2024agent) studies agent alignment in a multi-agent system, with evolutionary methods that go beyond Reinforcement Learning from Human Feedback (see Figure 30). In the context of agent alignment to norms, the approach is controlled: it does not permit the evolution of social norms to be disorderly or random, but it also does not intervene in each step of their evolution. The authors define the initial social norms and a desired direction of evolution. Agents with higher fitness (more norm-conforming) are more likely to reproduce, leading to the diffusion of their strategies, gradually stabilizing and forming new social norms.
Defining a complete, realistic, and complex virtual society is challenging. The purpose of the work is to study how, if such a virtual society existed, a system could further enable evolving intricate evolutionary behaviors of agents, and how this would lead to the emergence of new dynamics. The system provides a sandbox for investigating the safety of AI systems before they impact the real world.
A different approach is based on steels1995self’s naming game, implemented with agents powered by LLaMa 3 and Claude 3.5 (kouwenhoven2024searching; ashery2024dynamics; baronchelli2023shaping). They find that globally accepted conventions or norms can spontaneously arise from local interactions between communicating LLMs. The authors also demonstrate how strong collective biases can emerge during this process, even when individual agents appear to be unbiased, and how minority groups of committed LLMs can drive social change by establishing new social conventions that can overturn established behaviors.
The emergence of norms is studied at another level by horiguchi2024evolution. They explore the potential for LLM agents to spontaneously generate and adhere to normative strategies, building upon the foundational work of Axelrod’s metanorm games. Metanorms are norms enforcing the punishment of those who do not punish agents that are breaking norms (axelrod1986evolutionary). Controlling for personality traits vengefulness and boldness, they find that through dialogue, LLM agents can form complex social norms, metanorms, purely through natural language interaction. A related study evaluates the capability of LLMs to detect norm violations (he2024norm). Based on simulated data from 80 stories in a household context, they investigated whether 10 norms are violated, and found ChatGPT-4 being able for detect norm violations, with Mistral some distance behind.
qiu2024evaluating go a step beyond norms, and study the cultural and social awareness of LLM agents. They introduce CASA, a benchmark designed to assess LLM agents’ sensitivity to cultural and social norms across two web-based tasks: online shopping and social discussion forums. (CASA is based on WebArena (qi2024webrl).) Current LLMs perform significantly better in non-agent than in web-based agent environments, with agents achieving less than 10% awareness coverage and over 40% violation rates. However, using prompting and finetuning on specific datasets, cultural and social awareness can be improved.
Inspired by Society of Mind (minsky1988society), cooperation mechanisms are explored in zhang2023exploring’s agentic LLM simulation. This simulation consists of four unique societies of LLM agents, where each agent is characterized by a specific trait (easy-going or overconfident) and engages in cooperation with a distinct thinking pattern (debate or reflection). They find that LLM agents show human-like social behaviors, such as conformity and consensus reaching, mirroring foundational social psychology theories. Figure 31 shows societies with different types of agents.
The question whether groups of LLM agents can successfully engage in cross-national collaboration and debate is studied by (baltaji2024conformity). Multi agent discussions can support collective AI decisions that reflect diverse perspectives, although agents are susceptible to conformity due to perceived peer pressure. They can also lose track of their personas and opinions. Instructions that encourage debate increase the risk of errors.
4.3.3. Open-World Agents
An important driver of agentic LLM research is the problem of plateauing LLM performance due to limited training data. Open World multi-agent interaction aims to address this problem, generating new interaction data with multi agent simulations. Machine learning can learn no more complexity than what is present in the dataset (or environment). The idea of an open world-model is that it can create infinite datasets or environments, in which agents can continue to learn, to keep improving their intelligence. How should such unlimited challenges be created? The advent of LLMs has given a new impulse to this research question: LLMs are used to solve an LLM-generated problem. This idea is followed, for example, in the multi-agent finetuning approach (subramaniam2025multiagent).
Current agents are primarily created and tested in simplified synthetic environments, leading to a disconnect with real-world scenarios. zhou2023webarena build an environment, Webarena, that is more realistic and reproducible. WebArena is an environment with fully functional websites from four common domains: e-commerce, social forum discussions, collaborative software development, and content management.
Games are eminently suited as open-ended benchmarks for interactive behavior. Real-world tasks require handling intricate interactions, advanced spatial reasoning, long-term planning, and continuous exploration of new strategies. Balrog (paglieri2024balrog) incorporates reinforcement learning environments of varying levels of difficulty, ranging from tasks that are solvable by non-expert humans in seconds to challenging ones that may take years to master (such as the NetHack Learning Environment). They find that while current models achieve partial success in the easier games, they struggle significantly with more challenging tasks such as vision-based decision-making.
Progress in machine learning depends on benchmark availability. As models evolve, there is a need to create benchmarks that can measure progress on new generative capabilities (butt2024benchagents). BenchAgents decomposes the benchmark creation process into planning, generation, data verification, and evaluation, each of which is executed by an LLM agent. These agents interact with each other and utilize human-in-the-loop feedback to explicitly improve and flexibly control data diversity and quality. BenchAgents creates benchmarks to evaluate capabilities related to planning and constraint satisfaction.
AgentBoard provides an evaluation of the breadth of existing benchmarks (ma2024agentboard). Benchmarks should have task diversity. It is necessary to cover various agent tasks such as embodied, web, and tool tasks. Additionally, multi round interaction is important, to mimic realistic scenarios. Existing benchmarks typically adopt single-round tasks. Furthermore, agents should be evaluated in partially-observable environments, to test if they can actively explore their surroundings. Existing agent benchmarks fail to satisfy all of these criteria (ma2024agentboard).
4.4. Discussion
In this third part of the taxonomy, the focus was on agents that interact with other agents, both human and artificial. The goal is to understand social interaction, from interaction in conversations, social scenarios and dilemmas, to role-playing in duos and small teams, to large-scale open-ended emergent behavior at society level.
4.4.1. Interaction Studies
Over the past years, LLMs have provided us with new instances of human-machine interaction. Users across the globe have engaged in chat conversations seeking assistance with tasks in their professional or private lives. To engage in such interactions, LLMs rely on functions learned during training that we can recognize as social, including abilities for conversation, politeness and etiquette, handling of emotional and affective states, strategizing, and theory of mind. Such abilities form the basis not only for human-machine interaction, but also for LLM-driven machine-machine interactions, as we discuss next.
When interacting in multi-agent environments, agentic LLMs show varying levels of performance on games with strategic and coordinated behavior. Enhancing models with reasoning capacities boosts performance, as is evidenced by GPT-o1’s better overall performance and the positive effect of adding probabilistic graphs (xu2024magic). Pre-defined roles and interaction protocols (cooperative or adversarial) help structure the communication between LLM agents while improving task performance. Role-playing frameworks, AI feedback loops, and debate moderation suggests that carefully coordinating multiple LLMs can harness their collective intelligence and yield outcomes that surpass single model performance.
We have covered open-ended multi-LLM simulations without prior role assignment. These simulations give a new impulse to long-standing interests in the social sciences to model self-organizing behaviors, collective intelligence and the development of social conventions and norms. The scale of such simulations varies from a few interacting agents up to a million. Emergent behaviors are observed, such as coordination through norms and social structures that form spontaneously. In open-world approaches LLMs are used to create increasingly complex challenges and solutions for LLM agents.
4.4.2. In Depth: CAMEL and Generative Agents
We have surveyed the individual methods for social interaction. In order to dig deeper and highlight some of the issues in this area, we will now discuss two approaches, CAMEL, and Generative Agents, in more detail.
CAMEL
In Section 4.2.2 we saw how LLMs can be prompted to perform different roles, and work together in solving tasks. CAMEL (li2023camel) is a multi-agent system that has been designed to perform this role-playing-for-problem-solving task. Just like the assistants in Section 3.3, the goal is to solve a task, for example in medicine, finance, or computer programming (see also Figure 22). One approach is to write a single monolithic prompt for the LLM in which all the instructions to solve the task are specified—or so the author of the prompt hopes. To measure the success of this approach, a benchmark of test-cases can be created that the LLM has to solve.
The approach of CAMEL is different. CAMEL starts with the idea that specialized agents, with different prompts, may work better. Furthermore, the idea of CAMEL is that the different agents may work better together, and that new problem solving approaches may emerge, that were not present in the single monolithic prompt, or in the initial prompts of the individual agents. CAMEL is a multi-agent system that consists of two agents, an assistant and a user. The user has the domain knowledge, can provide a task specification, as well as feedback on intermediate deliverables that the assistant makes. The assistant has certain skills, for example Python coding, to write a program to solve the problem that the user agent specifies. Both agents are given a so-called inception prompt, an initial idea. They then further send messages (prompts) to each other as they work towards solving the task.
The CAMEL paper describes an experiment where a powerful LLM, GPT4, generates the initial prompt, and a cheaper LLM, GPT-3.5-turbo, executes the further steps. Ten different tasks have been simulated by a combination of 50 agents and 50 users, for a total of 25000 conversations (that were depth-limited to 40 messages).
In such a setup where agents work with agents, the training process may become unstable. CAMEL reports problems where the assistant simply repeats instructions (instead of answering them), problems with fake replies, empty replies, infinite loops of messages, and role reversal. A special (third) critic agent was introduced to ensure a constructive communication process between the agents.
The final experiments report success. The multi-agent system performed better than a single prompt, with performance on the HumanEval benchmark improving form 30% to around 50%. The CAMEL experiments showed that a multi-agent LLM system can be used for problem solving, and the authors showed how stable learning can be achieved when attention is paid to a constructive communication process.
Generative Agents
The architecture of CAMEL recognized two types of agents, assistant and user. The next two papers that we discuss, park2023generative; park2024generative, are about simulating a society with a larger number of agents, from 10-1000. The first paper (park2023generative) introduces the generative agents’ architecture, an architecture that was designed with the aim of generating believable proxies of human behavior. The architecture consists of a memory, a planning module, and a reflection module. In this design agents are given roles and operate in a Sims-like environment. An example is described where they decide to organize a Valentine’s day party (see also Figure 27, and the live web simulation).555See https://reverie.herokuapp.com/arXiv_Demo/ The focus of this work is on achieving believable social behavior, that is unscripted: behavior that emerges from the communications by the agents. We quote the paper: agents wake up, cook breakfast, and head to work; artists paint, while authors write; they form opinions, notice each other, and initiate conversations; they remember and reflect on days past as they plan the next day. The number of agents is larger than in CAMEL, arond 5-15, and the agent architecture is also more involved. Where in CAMEL the behavior was specified fully in-context, in the prompts that are exchanged, in Generative Agents there are additional external algorithms for memory, planning, and reflection. The paper further notes that: the new goals require architectures that manage constantly-growing memories as new interactions, conflicts, and events arise and fade over time while handling cascading social dynamics that unfold between multiple agents, and that success requires an approach that can retrieve relevant events and interactions over a long period, reflect on those memories to generalize and draw higher-level inferences, and apply that reasoning to create plans and reactions that make sense in the moment and in the longer-term of the agent’s behavior. The LLM that is used is GPT-3.5, the same as in CAMEL.
Whether the Generative Agents system generates believable behavior is evaluated with the help of 100 human evaluators. These were enlisted to rank believability of the communication patterns on four categories, by interviewing the agents to probe their ability to (1) remember past experiences, (2) plan future actions based on their experiences, (3) react appropriately to unexpected events, and (4) reflect on their performance to improve their future actions. The evaluation studied the results of four different agent architectures (full architecture, and no observation, no refection, no planning), and found that the full architecture performed best. They also found that generative agents remember with embellishments, and that reflection is required for synthesis of memories. The evaluation did find evidence of emerging communication, relationship building, and coordination. They also found evidence of erratic behavior, hallucination, and misclassification, such as agents that were trying to enter stores after closing time, not understanding the concept of closing time.
Simulating 1000 People
We will now turn to the second paper (park2024generative). Multi-agent simulations can be used to study different aspects of emergent individual and social human behavior. An important methodological challenge in prompt-based simulation studies is to determine how much of the behavior is scripted, and how much emerges. park2024generative focus in the second study on how realistic the behavior of synthetic agents can be. The study is based on their earlier work, with LLM agents that have memory and reflection. A group of 1052 human individuals were recruited who were asked to provide a two hour long interview. The interviews were standardized, administered by an AI interviewer. Next, LLM agents were trained on the audio interview, yielding 1052 different LLM agent profiles. The LLM is prompted to replicate individuals’ attitudes and behaviors and generate synthetic agents.
To validate the accuracy of the personality and behavior of these generative synthetic agents, the agents were tested by interviewing them. As a control group, the human subjects were also interviewed, again, two weeks after their initial interview, to control for natural variation between two interview sessions. The generative agents replicate participants’ responses on the General Social Survey 85% as accurately as participants replicate their own answers two weeks later, and perform comparably in predicting personality traits and outcomes in experimental replications. Subsequently, the synthetic agents have been made available for further experiments.
4.4.3. Emergent Collective Behavior
Emergent behavior, and especially emergent cooperation, is an important use case of agentic LLMs. It helps us understand our own behavior in our society, and allows the study of agent behavior in artificial conditions, in what-if scenarios. When do we benefit from more competition, when from more cooperation, and in what form? What happens when (fake) information disseminates? Or how do societies respond to extreme circumstances, such as a natural disaster?
As research on collective agent societies and emergent phenomena develops further, LLMs will exhibit more realistic behavior, new multi-agent infrastructures will be developed that allow more diverse types of interactions, and simulation studies will provide insight into social science questions. In particular, topics of interest are the influence of LLMs on democratic processes and cyber security, role playing, society of minds, theory of mind, curriculum learning, continuous learning, adversarial agents, and collaboration in the face of hierarchy.
Furthermore, as our understanding of the conditions conducive to emergence of cooperation grows, a focus on adaptive (social) intelligence may influence our views on the nature of intelligence and artificial (super)intelligence.
4.4.4. New Training Data
A final use case of this third part of the taxonomy is that new training data is generated by the interacting agents. Traditionally, LLMs are trained on a large static corpus of language data, that is taken from the internet, and ultimately based on human actions, using self-supervised learning methods. As illustrated by the cycle in Figure 1, interacting agentic LLMs enable self-learning, in the style of reinforcement learning. Reinforcement learning is used increasingly in LLM training, for example to train reasoning models by OpenAI (huang2024o1; wu2024comparative), and DeepSeek (guo2025deepseek). New reinforcement learning methods such as GRPO (shao2024deepseekmath) and RLVR (lambert2024tulu) already allow inference-time chains of thoughts to be used for finetuning.
In reinforcement learning, agents choose their own actions in the world, and are not limited by a pre-existing dataset. In principle, they can learn the full complexity of the world, including the effects of their own actions. A challenge in reinforcement learning is the instability caused by feedback loops. Past reinforcement learning successes have achieved stable training through diverse exploration and low learning rates, requiring large computational efforts (silver2016mastering; vinyals2019grandmaster; brown2019superhuman). Open-ended and open-world multi-agent simulation may provide an alternative way to create the necessary diversity for stable convergence.
5. General Discussion and Research Agenda
The interest in agentic LLMs is large, and many research efforts have appeared over a short period. We have reviewed the field, with an emphasis on the most recent works.
First, there is an interest from society in agentic LLMs. Agentic LLMs can assist us in our daily lives in many ways—from writing essays, booking flights, having pleasant and interesting conversations, folding our laundry, to making better medical diagnoses, performing better stock analyses, to support healthy lifestyle changes, to make sure we take our medicine, to assist us when we are less mobile. Tool use by assistants is enabled by technology from the first category of the taxonomy: reasoning LLMs, self-reflection and retrieval augmentation. Both reasoning and tool use support new forms of interaction, with human and artificial agents, further enhancing societal applicability of LLMs.
Second, there is an interest from science in agentic LLMs, inside the AI research community and beyond. Since LLM agents can now interact in natural language, agent behavior can be better understood, and multi-agent simulations can be made more realistic than before. Important questions in social and political science can be researched, such as in game theory (social dilemmas), social interaction (negotiation, theory of mind), and societal dynamics (cooperation, norms, extreme situations). Some of these goals are within reach, some have been realized already, and some are becoming a possibility. Also in research applications, agent interactions are enabled by the previous two categories: social behavior benefits from reasoning and self reflection, social actions are increasingly grounded, and information can be retrieved to further enhance understanding of social contexts.
Finally, agentic LLMs generate data that can augment inference-time behavior and on which models can be further pretrained and finetuned, improving LLMs beyond the plateau researchers have observed recently. Figure 1 illustrates this cycle of continuous improvement.
5.1. Research Agenda for Agentic LLM
Our survey has yielded interesting directions for a research agenda for agentic LLMs, which we will now discuss in more detail. Please refer to Table 4 for a summary of the agenda.
Training Data
The benefit from language corpuses that are used for pretraining of LLMs is said to be plateauing. To improve the performance of LLMs on language (and reasoning) tasks further, it is important to continue to acquire training data that is sufficiently novel and challenging from a token-prediction point of view. Such data can be generated by making LLMs interact with the world at inference time.
Currently, in most approaches that were discussed in Section 2, inference-time compute is only used to improve performance on reasoning benchmarks. In most early Chain of Thought approaches the generated data is not used after the answer has been calculated. In other approaches—such as Say Can, Inner Monologue, and Vision-Language-Action models—data that is generated at inference time is used for augmentation of the finetuning dataset, creating an inference time-finetuning feedback loop, so that the model’s parameters are trained from its own earlier reasoning.
Such feedback loops are common in reinforcement learning, where agents act and receive feedback from their environment. In games, a self-learning loop can be created (plaat2022deep). In AlphaGo Zero this approach yielded good results, although at the cost of careful tuning of hyperparameters and algorithms, to ensure sustained convergence of the learning process (silver2017mastering). Similar results were achieved in other challenging games, such as StarCraft (vinyals2019grandmaster), Stratego (perolat2022mastering), DOTA 2 (berner2019dota), Diplomacy (meta2022human), and Poker (brown2019superhuman).
More formally, in the traditional self-supervised view a model is trained to predict label from input variable in dataset ; in reinforcement learning an agent’s policy is trained with reward to perform action to change state of its environment . In agentic LLMs, both views are joined. Agentic LLMs use a language model as the policy to determine the agent’s next action (see Figure 32).
Actions can be used to retrieve information, to split a larger problem into smaller parts, to run a tool, to use memory to reflect on its own actions, to suggest stock trades, to book travel tickets, or to interact with other agents working towards a common goal.
The approach that worked well in games of strategy is now also successfully used in robotics, in the creation of Vision-Language-Action models (black2024pi_0; brohan2023rt). VLAs that are trained on self-generated action sequences show zero-shot generalization results in domestic tasks (kitchen tasks, folding laundry) that had not been achieved by other machine learning methods. Recently, reasoning models—such as DeepSeek (guo2025deepseek) and Kimi (team2025kimi)—are also being trained with reinforcement learning. Popular methods for finetuning for mathematics and coding tasks are GRPO (shao2024deepseekmath) and RLVR (lambert2024tulu). Other uses of agents for finetuning are reported by subramaniam2025multiagent.
Reuse of inference time results for finetuning and pretraining closes the learning loop (see Figure 1), and is the first item for the agenda for further research. It is interesting to see how the reinforcement learning methods that worked well for games of strategy and certain finetuning tasks, are being translated to work in LLMs that act in the real world.
VLAs integrate multiple modalities: language, visual information, and actions. Further modalities are speech, other audio signals, and videos. Electrical signals, such as brain or muscle activation, can also provide valuable inputs for the models to learn from.
Hallucination and Stable Behavior
A challenge for the virtuous autocurriculum cycle is that LLMs hallucinate, and in multi-step reasoning errors can easily accumulate. LLM answers may look good, but be factually wrong. Reasoning chains may be unfaithful, giving good answers for the wrong reason, and wrong answers when least expected. Especially when such dubious results are used to further train the LLM, this training may diverge and model collapse may occur. In social simulations, emerging behavior patterns, such as cooperation, fairness, trust or norms, may collapse. Therefore, in multi-step reasoning, self verification and self consistency methods were developed to address error accumulation. In reinforcement learning, exploration and diversity are important methods to ensure good coverage of the state space. In social simulations and gaming, open world models and open-ended behavior are being used to stimulate exploration and diversity. Such models can provide suitable environments for automated generation of training curricula.
Faithfulness for Chain of Thought is studied by lyu2023faithful; lanham2023measuring; turpin2024language. Mechanistic interpretability can provide ways to look inside the LLM, to better understand if the model follows the reasoning steps that we expect it to take (nanda2023progress; bereska2024mechanistic; ferrando2024primer; chen2025does). The conditions that influence stability of emergent behavior (cooperation, fairness, trust) may be studied further.
For agentic LLMs that learn from their own results, other methods must be developed, and hallucination features prominently on the research agenda for agentic LLMs, with mechanistic interpretability and open world models as important items.
Agent Behavior at Scale
Studies of emergent behavior need realistic agent behavior, and we expect more research to be performed to improve agent behavior, for example by closely modeling human behavior in generative agents (park2024generative). Some behavior patterns in multi-agent simulations only emerge at scale, as studies with specialized agent infrastructures have shown (park2023generative; yang2024oasis; wu2023autogen). However, the number of LLM agents that can be simulated reliably is often limited. Although open-ended simulation show improved scalability, we believe that more research into scaling of simulations with complex agents is necessary.
Related to the challenge of scale is the cost of training LLMs. Pretraining and finetuning an LLM is expensive. Knowledge distillation is a popular method to extract essential knowledge and behavior from a large model into a small model, at lower computational cost (xu2024survey). Experiments have shown that reasoning steps can be distilled from large to smaller language models (gu2023minillm; li2023symbolic; muennighoff2025s1). Knowledge distillation in LLM agents in an important item for our research agenda.
Another aspect of agentic LLM research is the study of emergent behavior, of cooperation and trust in agentic societies. The debate on artificial super-intelligence is fueled, in part, by the growing performance of individual LLMs, which is an important aspect of agentic LLM research. Studies of emergent agent behavior at scale may show us when cooperation and trust emerge, may influence our view on the nature of intelligence, and may thus influence the discussion on artificial super-intelligence and the future of society. Furthermore, the world around us is organized in groups in which power hierarchies are prevalent. Many multi agent simulations assume a flat power hierarchy. Multi agent simulations should also go beyond equality.
Self Reflection
Self reflection mechanisms are used in advanced prompt-improvement algorithms. Hand-writing external prompt management algorithms may be error prone and brittle. An alternative is to let the LLM perform the self reflection and step-by-step management internally, as in the original Chain of Thought (implicit reasoning (li2025implicit)).
DeepSeek R1 (guo2025deepseek) is a reasoning model that is trained (finetuned) by the GRPO reinforcement learning method (shao2024deepseekmath). The model is trained on its own reasoning results, and was found to self-reflectively reason over its own results, identifying effective reasoning patterns implicitly. schultz2024mastering train a model on search sequences (gandhi2024stream) in games such as chess, and VLAs are trained on action sequences (kim2024openvla). These works shows that, in addition to implicit step-by-step reasoning, implicit search is possible. An open question is whether LLMs can perform self reflection internally.
By adding external state to an LLM, we enable reasoning and a form of self reflection, which is a rudimentary form of metacognition (thinking about thinking). LLMs that reflect on their own behavior raise visions of true artificial intelligence. If LLMs can self-reflect, can they exhibit metacognition (wang2023metacognitive; didolkar2024metacognitive)? Self reflection by LLMs is another item for the research agenda.
When we add outside state to the input prompts, the input to the LLM will differ based on the history, and so will the answers of the LLM. Differences in memory may be preceived as a personality of the LLMs by its users. The question if LLMs with outside memory exhibit a personality is a topic for future research.
Self reflective methods are being used to create agents to perform scientific discovery (eger2025transforming). How these agents will influence, and possibly improve, the process of scientific discovery is an exciting area of research.
Safety
Safety is a crucial issue in LLMs that act in the world. The problem is studied, but far from being solved (brunke2022safe; andriushchenko2024agentharm; samvelyan2024rainbow). Actions by assistants and robots in the real world have real world consequences. When a financial trading assistant hallucinates, or when a self driving robot makes a wrong inference, questions on responsibility and liability should be addressed. More legal and ethical questions arise, for example, on privacy and fairness, and, possibly, concerning the rights of algorithmic entities (harris2021moral; bengio2025illusions).
The application areas for the assistants in this survey—shopping, medical diagnosis, finance—are narrow. The narrower the application domain, the better the answers.
Ensuring the safety of agentic LLMs requires moving beyond prompt-based defenses toward integrated, multi-layer safeguards. Key directions include embedding explicit safety constraints within the agent’s reasoning and planning pipeline and employing continual adversarial training and automated red-teaming to enhance robustness against manipulation. Further priorities are developing mechanisms for self-regulation and risk awareness, enabling agents to detect and avoid unsafe actions, and establishing rigorous, standardized safety benchmarks such as AgentHarm (andriushchenko2024agentharm). Together, these measures outline a roadmap for accountable and trustworthy deployment of agentic LLMs in high-stakes domains.
Clearly, many safety, ethics and trust issues will have to be addressed before the full breadth of the possibilities of agentic LLMs can be enjoyed. Safety and ethics will become important topics on the research agenda of agentic LLMs, deserving their own surveys and books (gan2024navigating; jiao2025navigating; raza2025responsible).
| Topic | Challenge |
| Training Data | Finetune with inference time reasoning data |
| Convergent/stable reinforcement learning | |
| VLA, Multimodal signals, such as speech | |
| Hallucination | Use Self Verification |
| Use Mechanistic Interpretability | |
| Use Open Ended/Open World Models for exploration | |
| Agent Behavior | Scalable simulation infrastructure, role playing |
| Distill reasoning to small models | |
| Models of agent and human behavior, emergent behavior, future of society | |
| Self reflection | In-model self reflection and metareasoning |
| Metacognition, personality | |
| Automated Scientific Discovery | |
| Safety | Assistants: Responsibility, liability |
| Privacy, fairness of data | |
| Wider application areas for assistants |
5.2. Conclusion
There is a large research activity on agentic LLMs. Already, robots show impressive generalization results, and so do assistants in medical diagnosis, financial market advising, and scientific research. Work processes in these—and other—fields may well be affected by agentic LLM assistants in the near future.
The agentic LLMs in this survey have (1) reasoning capabilities, (2) an interface to the outside world in order to act, and (3) a social environment with other agents with which to interact. The categories of this taxonomy complement each other. At the basis is the reasoning technology of category 1. Robotic interaction and tool-use build on grounded retrieval augmentation, social interaction (such as theory of mind) builds on self reflection, and all categories benefit from reasoning and self-verification. Closing the cycle, the acting and interacting categories generate training data for further pretraining and finetuning LLMs, beyond plateauing traditional datasets (Figure 1). The impressive generalization capabilities of Vision-Language-Action models are testament to the power of this approach.
The reasoning paradigm connects to works in human cognition, and some papers anthropomorphize LLM computations in Kahneman’s terms of System 1 thinking (fast, associative) and System 2 thinking (slow, deliberative). Works on reasoning focus on the intelligence of single LLMs. This individualistic view also gives rise to discussions about superintelligence, some utopian, some not.
The agentic paradigm enables two elements of machine learning that are new for LLMs. First, in reinforcement learning, agents self reflect and choose their own actions, and learn from the feedback of the world in which they operate. Second, no dataset is needed, nor is learning limited by the complexity of a dataset, it is only limited by the complexity of the world around the agent. The agent paradigm creates a more challenging training setting, allowing agentic LLMs to keep improving themselves.
The multi-agent paradigm studies agent-agent societies. The focus is on emergent behaviors such as egoism/altruism, competition/collaboration, and (dis)trust. Social cognitive development and the emergence of collective intelligence are also studied in this field. Connecting back to the reasoning paradigm, the collaboratieve view of multi-agent studies may inform discussions about (super)intelligence, teaching us about emerging social behavior of LLM-agents.
Acknowledgements
We thank Joost Broekens, Suzan Verberne, Annie Wong, Thomas Bäck, Zhaochun Ren, Thomas Moerland, Michiel van der Meer, Bram Renting, Frank van Harmelen, Tessa Verhoef, and Rob van Nieuwpoort for extensive and fruitful discussions. We thank the anonymous reviewers for valuable suggestions that have improved the article.
6. Reproducibility Checklist for JAIR
Select the answers that apply to your research – one per item.
All articles:
All claims investigated in this work are clearly stated. [yes]
Clear explanations are given how the work reported substantiates the claims. [yes]
Limitations or technical assumptions are stated clearly and explicitly. [yes]
Conceptual outlines and/or pseudo-code descriptions of the AI methods introduced in this work are provided, and important implementation details are discussed. [yes]
Motivation is provided for all design choices, including algorithms, implementation choices, parameters, data sets and experimental protocols beyond metrics. [yes]
Articles containing theoretical contributions:
Does this paper make theoretical contributions? [no]
If yes, please complete the list below.
All assumptions and restrictions are stated clearly and formally. [yes/partially/no]
All novel claims are stated formally (e.g., in theorem statements). [yes/partially/no]
Proofs of all non-trivial claims are provided in sufficient detail to permit verification by readers with a reasonable degree of expertise (e.g., that expected from a PhD candidate in the same area of AI). [yes/partially/no]
Complex formalism, such as definitions or proofs, is motivated and explained clearly. [yes/partially/no]
The use of mathematical notation and formalism serves the purpose of enhancing clarity and precision; gratuitous use of mathematical formalism (i.e., use that does not enhance clarity or precision) is avoided. [yes/partially/no]
Appropriate citations are given for all non-trivial theoretical tools and techniques. [yes/partially/no]
Articles reporting on computational experiments:
Does this paper include computational experiments? [no]
If yes, please complete the list below.
All source code required for conducting experiments is included in an online appendix or will be made publicly available upon publication of the paper. The online appendix follows best practices for source code readability and documentation as well as for long-term accessibility. [yes/partially/no]
The source code comes with a license that allows free usage for reproducibility purposes. [yes/partially/no]
The source code comes with a license that allows free usage for research purposes in general. [yes/partially/no]
Raw, unaggregated data from all experiments is included in an online appendix or will be made publicly available upon publication of the paper. The online appendix follows best practices for long-term accessibility. [yes/partially/no]
The unaggregated data comes with a license that allows free usage for reproducibility purposes. [yes/partially/no]
The unaggregated data comes with a license that allows free usage for research purposes in general. [yes/partially/no]
If an algorithm depends on randomness, then the method used for generating random numbers and for setting seeds is described in a way sufficient to allow replication of results. [yes/partially/no/NA]
The execution environment for experiments, the computing infrastructure (hardware and software) used for running them, is described, including GPU/CPU makes and models; amount of memory (cache and RAM); make and version of operating system; names and versions of relevant software libraries and frameworks. [yes/partially/no]
The evaluation metrics used in experiments are clearly explained and their choice is explicitly motivated. [yes/partially/no]
The number of algorithm runs used to compute each result is reported. [yes/no]
Reported results have not been “cherry-picked” by silently ignoring unsuccessful or unsatisfactory experiments. [yes/partially/no]
Analysis of results goes beyond single-dimensional summaries of performance (e.g., average, median) to include measures of variation, confidence, or other distributional information. [yes/no]
All (hyper-) parameter settings for the algorithms/methods used in experiments have been reported, along with the rationale or method for determining them. [yes/partially/no/NA]
The number and range of (hyper-) parameter settings explored prior to conducting final experiments have been indicated, along with the effort spent on (hyper-) parameter optimisation. [yes/partially/no/NA]
Appropriately chosen statistical hypothesis tests are used to establish statistical significance in the presence of noise effects. [yes/partially/no/NA]
Articles using data sets:
Does this work rely on one or more data sets (possibly obtained from a benchmark generator or similar software artifact)? [no]
If yes, please complete the list below.
All newly introduced data sets are included in an online appendix or will be made publicly available upon publication of the paper. The online appendix follows best practices for long-term accessibility with a license that allows free usage for research purposes. [yes/partially/no/NA]
The newly introduced data set comes with a license that allows free usage for reproducibility purposes. [yes/partially/no]
The newly introduced data set comes with a license that allows free usage for research purposes in general. [yes/partially/no]
All data sets drawn from the literature or other public sources (potentially including authors’ own previously published work) are accompanied by appropriate citations. [yes/no/NA]
All data sets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. [yes/partially/no/NA]
All new data sets and data sets that are not publicly available are described in detail, including relevant statistics, the data collection process and annotation process if relevant. [yes/partially/no/NA]
All methods used for preprocessing, augmenting, batching or splitting data sets (e.g., in the context of hold-out or cross-validation) are described in detail. [yes/partially/no/NA]